### v1

# Frontiers in Intelligent Colonoscopy

Abstract
Colonoscopy is currently one of the most sensitive screening methods for colorectal cancer. This study investigates the frontiers of intelligent colonoscopy techniques and their prospective implications for multimodal medical applications. With this goal, we begin by assessing the current data-centric and model-centric landscapes through four tasks for colonoscopic scene perception, including classification, detection, segmentation, and vision-language understanding. This assessment enables us to identify domain-specific challenges and reveals that multimodal research in colonoscopy remains open for further exploration. To embrace the coming multimodal era, we establish three foundational initiatives: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-designed multimodal language model ColonGPT, and a multimodal benchmark. To facilitate ongoing monitoring of this rapidly evolving field, we provide a public website for the latest updates: https://github.com/ai4colonoscopy/IntelliScope.
Index Terms:
Colonoscopy, Survey, Vision-language, Multimodal Language Model, Medical Image, Abdomen, Healthcare AI.
1 Introduction
Despite declining colorectal cancer (CRC) rates in high-income countries, it remains the third most diagnosed cancer worldwide and is increasing in developing countries [1]. Colonoscopy, as an efficient method for CRC screening, utilises a flexible camera-equipped tube to visually examine the colon’s interior. As illustrated in Fig. (1-a), this clinical procedure also facilitates intervention with specialised instruments such as snares, forceps, and cautery devices to remove precancerous growths, such as serrated and adenomatous polyps. A recent study [2] indicates that incorporating artificial intelligence (AI) into colonoscopy reduces the miss rate of colorectal neoplasia by approximately 50% compared to traditional methods. This success motivates us to investigate the frontiers in intelligent colonoscopy.
Colonoscopy, an endoscopic optical imaging technique, usually presents visual patterns (e.g., non-uniform illumination, homogeneity) that differ from those of general-purpose imaging data, e.g., ImageNet [3], due to the complex and folded anatomy of the colon. This suggests that special methods are needed to interpret the colonoscopic data. In response, we begin with an investigation of the latest intelligent techniques for colonoscopy, assessing the current landscape to sort out domain-unique challenges and underexplored areas. Our analysis reveals that multimodal research in colonoscopy remains largely untapped. To bridge this gap, we contribute three efforts to the community, as illustrated in Fig. (1-b).
Figure 1: Introductory diagram. We depict (a) the anatomy of the large intestine (colon) within the digestive tract, the polypectomy procedure during colonoscopy examination, and the components of a colonoscope. The bottom figure (b) summarises three highlights of this study.
Figure 2: Colonscopic scene perception from visual to multimodal perspectives. In clinical practice, purely visual tasks, including (a) classification, (b) detection, and (c) segmentation, are applied to identify targets of interest such as polyps and instruments. (d) Multimodal applications improve colonoscopy procedures by performing interactive, user-driven tasks aligned with clinical needs. The chatbot provides personalised advice, automated reporting, and streamline procedural workflows.
Contribution. (a) We investigate the latest research progress in four colonoscopic scene perception tasks (refer to Fig. 2) from both data-centric and model-centric perspectives. Our investigation summarises key features of 63 datasets and 137 representative deep techniques published since 2015. Additionally, we highlight emerging trends and opportunities for future study. (b) We introduce ColonINST, a pioneering instruction tuning dataset tailored for multimodal research, aimed at instructing models to execute user-driven tasks interactively. Assembled from 19 publicly available sources, the ColonINST dataset contains 303,001 colonoscopy images across 62 sub-categories, reflecting diverse scenarios encountered in colonoscopy procedures. We expand these visual samples in two aspects. First, we leverage the multimodal AI chatbot, GPT-4V [4], to generate 128,620 medical captions. Second, we restructure 450,724 human-machine conversations for multimodal adaptation. (c) Leveraging the instruction tuning data, we build a multimodal language model, ColonGPT  to assist endoscopists through interactive dialogues. To ensure reproducibility for average community users, we implement ColonGPT in a resource-friendly way, using a 0.4B-parameter visual encoder SigLIP-SO [5] and a 1.3B-parameter lightweight language model Phi1.5 [6]. Unlike previous vision-language (VL) linking methods [7, 8, 9] that employ multilayer perceptrons to handle equally all tokens from the visual encoder, we propose a multigranularity adapter to selectively sample visual tokens according to their significance. This strategy reduces the visual tokens to $\sim$34of the original number without compromising performance, securing the top spot in our newly-created multimodal benchmark across three tasks. Importantly, our ColonGPT can be trained within five hours on four A100-40GB GPUs, facilitating rapid proof-of-concept development for subsequent research.
Scope. This study differs from previous works in several aspects. Earlier surveys on traditional [10] and deep learning [11, 12, 13, 14] methods conducted before 2020 are now out of date. Although a recent study [15] explores various applications of colonoscopy, such as quality analysis and abnormality detection, it lacks numerical validation. Other benchmarks [16, 17, 18] are limited to specific narrow research subfields. By contrast, we delve into four tasks for colonoscopic scene perception and evaluate their current state to sort out key challenges and under-researched areas. Importantly, our vision goes beyond by laying the foundations for the coming era, a multimodal world. To embrace this era, we further undertake three initiatives: a multimodal instruction tuning dataset, a multimodal language model, and a multimodal benchmark for the community.
Organisation. The remaining sections are structured as such: §2 provides a historical background and discusses the domain-unique challenges. §3 investigates 63 colonoscopy related datasets, followed by a survey of 137 deep models in §4. In §5, we introduce three initiatives towards the multimodal era: the creation of ColonINST, the technical details of ColonGPT, and a comparative multimodal benchmark along with ablative analyses. Finally, this paper is concluded in §6.
TABLE I: Data statistics for colonoscopy datasets. The columns include: number of images (#IMG) and videos (#VID), classification tag (Cls), bounding box (Bbx), segmentation mask (Seg), text (Tx). The categories not related to colonoscopy, such as stomach and esophagitis, are marked in grey.
Dataset Publication #IMG #VID Cls Bbx Seg Tx Number of categories (#C) $\rightarrow$ Category names URL
CVC-ColonDB [19] PR’12 300 15 #C1 $\rightarrow$ polyp Link
ETIS-Larib [20] CARS’14 196 - #C2 $\rightarrow$ polyp, non-polyp Link
CVC-ClinicDB [21] CMIG’15 612 31 #C1 $\rightarrow$ polyp Link
ASU-Mayo [22] TMI’15 36,458 38 #C2 $\rightarrow$ polyp, non-polyp Link
Ye et al.[23] MedIA’16 7,894 10 #C2 $\rightarrow$ polyp, non-instance Link
Deeba et al.[24] IJCNN’16 100 - #C2 $\rightarrow$ bleeding, non-bleeding -
CU-ColonDB [25] JBHI’16 1,930 - #C3 $\rightarrow$ hyperplasia polyps, adenomatous polyps, non-polyp -
ColonoscopicDS [26] TMI’16 - 76 #C3 $\rightarrow$ serrated adenomas, hyperplastic lesions, adenoma Link
CVC-ClinicVideoDB [27] MICCAIw’17 10,924 18 #C2 $\rightarrow$ polyp, non-polyp Link
Kvasir [28] MMSys’17 8,000 - #C8 $\rightarrow$ cecum, polyps, ulcerative colitis, dyed and lifted polyp, dyed resection margins, Z-line, pylorus, esophagitis Link
Nerthus [29] MMSys’17 5,525 21 #C4 $\rightarrow$ BBPS (Boston-Bowel-Preparation-Scale) 0/1/2/3 Link
EndoSceneStill [30] JHE’17 912 44 #C1 $\rightarrow$ polyp Link
KID1 [31] EIO’17 137 - #C10 $\rightarrow$ angiectasias, ulcers, stenoses, villous edema, nodular lymphangiectasias, chylous cysts, polyps, aphthae, normal/no pathology, intraluminal hemorrhage Link
KID2 [31] EIO’17 2,371 47 #C4 $\rightarrow$ vascular anomalies, polypoid anomalies, inflammatory anomalies, normal images Link
NBIPolyp-UCdb [32] BSPC’19 86 11 #C2 $\rightarrow$ adenomas, hyperplastic Link
WLPolyp-UCdb [33] EIO’19 3,040 42 #C2 $\rightarrow$ polyp, normal mucosa Link
ASEI [34] MM’19 4,470 - #C4 $\rightarrow$ dyed-lifted-polyps, dyed-resection-margins, instruments, polyp Link
Cho et al.[35] PeerJ’19 328,927 112 #C1 $\rightarrow$ cecum Link
EAD2019 [36] arXiv’ 19 2,342 - #C7 $\rightarrow$ imaging artefacts, contrast, specularity, instrument, bubbles, motion blur, saturation Link
Liu et al.[37] ISBI’20 14,317 18 #C2 $\rightarrow$ polyp, non-polyp -
Kvasir-SEG [38] MMM’20 1,000 - #C1 $\rightarrow$ polyp Link
PICCOLO [39] ApplSci’20 3,433 39 #C17 $\rightarrow$ Paris classification (protruded lesions: 0-Ip/0-Ips/0-Is, elevated lesions: 0-IIa/0-IIa+c, flat lesions: 0-IIb), NICE classification (type 1/2/3), Diagnosis (adenocarcinoma/adenoma/hyperplasia), Histological stratification (high grade dysplasia/hyperplasia/invasive adenocarcinoma/low grade dysplasia/no dysplasia) Link
EDD2020 [40] arXiv’20 386 - #C5 $\rightarrow$ suspicious area, high-grade dysplasia, adenocarcinoma, polyp, normal dysplastic Barrett’s oesophagus Link
CAD-CAP [41] EIO’20 25,124 1,686 #C4 $\rightarrow$ vascular lesions, fresh blood, ulcero-inflammatory lesions, normal images -
ACP-ColonDB530 [42] NPJDM’20 221,976 - #C13 $\rightarrow$ adenomatous polyp, hyperplastic polyp, other polyp, bleeding, IC valve, instrument, artefact, normal colon structure, bubble, inside colon background, stool, lumen, outside colon background -
HyperKvasir [43] SData’20 110,079 374 #C23 $\rightarrow$ cecum, retroflex rectum, BBPS 0-1/2-3, ulcerative colitis grade 1/2/3/0-1/1-2/2-3, polyps, dyed lifted polyps, dyed resection margins, hemorrhoids, Barrett’s, terminal ileum, Z-line, esophagitis grade A, esophagitis grade B-D, pylorus, retroflex stomach, Barrett’s (short-segment), impacted stool Link
WCE-Polyp [44] TMI’20 541 - #C1 $\rightarrow$ polyp Link
EAD2020 [45] MedIA’21 2,531 - #C8 $\rightarrow$ specularity, bubbles, saturation, contrast, blood, instrument, blur, imaging artefacts Link
BKAI-Small [46] ISVC’21 1,200 - #C3 $\rightarrow$ non-neoplastic polyp, neoplastic polyp, background Link
BKAI-Large [46] ISVC’21 7,466 - #C4 $\rightarrow$ non-neoplastic polyp, neoplastic polyp, undefined polyp, background Link
CPC-Paired [47] MICCAI’21 681 - #C2 $\rightarrow$ hyperplastic polyp, adenoma Link
LDPolyVideo [48] MICCAI’21 901,666 263 #C2 $\rightarrow$ polyp, non-polyp Link
Celik et al.[49] MICCAI’21 2,224 - #C2 $\rightarrow$ polyps, Barrett’s esophagus Link
Kvasir-Instrument [50] MMM’21 590 - #C1 $\rightarrow$ GI procedure tools (e.g., snares, balloons, and biopsy forceps) Link
CP-CHILD [51] BMCMI’21 9,500 - #C2 $\rightarrow$ colonic polyp, normal or other pathological images Link
CROHN-IPI [52] EIO’21 3,498 - #C7 $\rightarrow$ erythema, edema, aphthoid ulceration, ulceration (3–10mm, $>$10mm), stenosis, non-pathological Link
C-E Crohn’s Disease [53] FMOLB’21 467 164 #C1 $\rightarrow$ Crohn’s lesions -
SUN-database [54] GIE’21 159,232 113 #C7 $\rightarrow$ hyperplastic polyp, low grade adenoma, high-grade adenoma, traditional serrated adenoma, invasive carcinoma, sessile serrated lesion, negative Link
Kvasir-Sessile [55] JBHI’21 196 - #C1 $\rightarrow$ polyp ($<$10mm) Link
Kvasir-Capsule [56] SData’21 4,741,504 117 #C14 $\rightarrow$ polyp, Ileocecal valve, lymphangiectasia, erythema, angiectasia, foreign body, erosion, ulcer, blood (fresh), blood (hematin), normal clean mucosa, reduced mucosal view, pylorus, ampulla of Vater Link
KUMC [57] PONE’21 37,899 155 #C2 $\rightarrow$ hyperplastic polyps, adenomatous polyps Link
ERS* [58] arXiv’22 1,354,667 1,520 #C27 $\rightarrow$ ulcerative colitis (active/quiescent), stricture (postoperative/inflammatory/malignant), polyp, melanosis, diverticulosis, fistula, crohnsdisease (active/quiescent), lipoma, proctitis, hemorrhoids, submucosal tumor, solitary ulcer, bleeding of unknown origin, ileitis, diverticulitis, colitis: ischemic, colorectal cancer, angiodysplasia, rectal ulcer, foreign body, polyposis syndrome, postoperative appearance, parasites Link
Tian et al.[59] MICCAI’22 807,069 253 #C2 $\rightarrow$ polyp, non-polyp Link
WCE-CCDD [60] BSPC’22 6,000 - #C4 $\rightarrow$ ulcer, polyps, normal, esophagitis Link
PolypGen2.0 [61] ISBIw’22 3,446 46 #C2 $\rightarrow$ serrated, adenomas Link
SUN-SEG [62] MIR’22 159,232 1,013 #C7 $\rightarrow$ hyperplastic polyp, low grade adenoma, high-grade adenoma, traditional serrated adenoma, invasive carcinoma, sessile serrated lesion, negative Link
SinGAN-Seg [63] PONE’22 10,000 - #C1 $\rightarrow$ polyp Link
ENDOTEST [64] SJG’22 253,754 58 #C2 $\rightarrow$ polyp, non-polyp Link
MEDVQA-GI [65] CLEF’23 3,949 - #C2 $\rightarrow$ polyp, surgical equipment Link
GastroVision [66] ICMLw’23 8,000 - #C27 $\rightarrow$ accessory tools, angiectasia, blood in lumen, cecum, colon diverticula, resection margins, colorectal cancer, dyed-lifted-polyps, erythema, ulcer, dyed-resection-margins, retroflex rectum, mucosal inflammation large bowel, resected polyps, colon polyps, lleocecal valve, normal mucosa and vascular pattern in the large bowel, esophagitis, Barrett’s esophagus, duodenal bulb, esophageal varices, gastric polyps, gastroesophageal junction normal z-line, normal esophagus, normal stomach, pylorus, small bowel terminal ileum Link
W-Polyp [67] CVPR’23 1,450 - #C1 $\rightarrow$ polyp Link
LIMUC [68] IBD’23 11,276 - #C4 $\rightarrow$ Mayo endoscopic score (MES) 0/1/2/3 Link
PS-NBI2K [16] JBHI’23 2,000 - #C1 $\rightarrow$ polyp Link
PolypGen [69] SData’23 8,037 23 #C2 $\rightarrow$ polyp, negative Link
MedFMC** [70] SData’23 22,349 - #C5 $\rightarrow$ ulcer, erosion, polyp, tumor, and non-instance Link
GB-WCE Dataset [71] MD’23 226 - #C2 $\rightarrow$ bleeding or lesions, normal Link
REAL-Colon [72] SData’24 2,757,723 60 #C2 $\rightarrow$ polyp, negative Link
Xu et al.[73] TMI’24 251 - #C4 $\rightarrow$ Mayo endoscopic score (MES) 0/1/2/3 Link
Kvasir-VQA [74] MMw’24 6,500 - #C5 $\rightarrow$ polyps, ulcerative colitis, instrument, normal, esophagitis Link
CapsuleVision2024 [75] CVIP’24 58,124 - #C10 $\rightarrow$ angioectasia, bleeding, erosion, erythema, foreign body, lymphangiectasia, polyp, ulcer, worms, normal Link
COLON [76] arXiv’24 $\sim$430,000 30 #C3 $\rightarrow$ adenoma, hyperplastic, non-pathological case -
WCEBleedGen [77] arXiv’24 2,618 - #C2 $\rightarrow$ bleeding, non-bleeding Link
PolypDB [78] arXiv’24 3,934 - #C1 $\rightarrow$ polyp (multiple imaging modalities and multiple medical centers) Link
  • *NOTE – The ERS dataset [58] includes 99 annotated categories in total. For the sake of brevity, we list only 27 colon-related categories within ERS.
  • **NOTE – The MedFMC dataset [70] comprises 23,349 medical images across five modalities. This table only enumerates the categories specific to the endoscopic modality.
2 Background
2.1 Origin and evolution
The history of colonoscopy has two key milestones. In 1968, gastrointestinal surgeons Hiromi Shinya and William Wolff found a link between colonic polyps and intestinal tumours, but they lacked equipment to examine them. In 1969, they discovered Corning Incorporated’s optical fibres and collaborated with Olympus to create the fiberoptic colonoscope, a groundbreaking device to examine the colon and remove polyps using wire loops. The second milestone came in 1983 with the introduction of the electronic colonoscope [79], which allows visualisation of the colon on a screen and polyp removal using a polypectomy snare, enhancing detection rates and reducing bleeding. The 21st century brings the AI era, where computer-aided diagnosis systems provide greater precision and reliability in procedures [80]. This study explores the transformative impact of intelligent techniques for colonoscopy, which is a type of endoscopy [81], while other related techniques such as laparoscopy [82] are summarised in our appendix.
2.2 Intrinsic traits and domain-unique challenges
We summarise five unique challenges associated with colonoscopic vision tasks, primarily caused by procedural aspects and imaging conditions during a colonoscopy. (a) Non-linear camera ego-motion. Procedural constraints force the camera (i.e., colonoscope) to actively move in a non-linear and unpredictable manner, challenging ego-motion compensation [83] and causing motion blur [62]. (b) Presence of medical instruments. The colonoscopy procedure often includes instruments such as scopes, guidewires, and snares, which should be distinguished properly from anatomical structures [84] for efficient analysis. (c) Limited observable field. The intricate folds and blind spots within the colon restrict the visible area in colonoscopy data. This requires algorithms capable of extracting relevant information from limited visual landscapes [85]. (d) Non-uniform illumination. The mucosal surface of the colon, prone to wetness and sheen, results in highly variable and diffuse illumination with complex reflections such as non-Lambertian reflections and interreflections. Traditional lighting-based algorithms struggle under these conditions [86]. (e) Variability in tissue appearance. Mucosal textures and colours vary considerably due to constant movement, disease states, anatomical differences, and instrument effects. Furthermore, benign polyps or lesions usually have weak or homogeneous boundaries [87], making them blend into surrounding tissues and difficult to detect. These issues require a robust response from AI models to inherent morphological and colour fluctuations.
3 Revisiting Colonoscopy Data
3.1 Medical data for colonoscopy
Tab. I presents our investigation of 63 datasets with their essential statistics for four tasks related to colonoscopic scene perception. We search for them using queries such as “colonoscopy dataset/benchmark” and “gastrointestinal disease dataset”. They consist of images or videos related to the human colon. In particular, some datasets also include images of other organs, such as the pylorus in [28] or the stomach in [66]. Next, we review these selected datasets according to their different task objectives.
$\bullet$ Classification datasets have been widely used for varied purposes, such as colon disease classification in images [52, 60, 68, 73, 28, 40, 66, 43, 31, 58, 75]/videos [54, 62, 56], polyp identification [27, 51, 33, 48, 37, 41, 39], fine-grained polyp classification [25, 47, 26], bleeding condition [24, 77, 71], anomaly recognition [59], cecum recognition [35], and pre-operative assessment [29].
$\bullet$ Detection datasets provide both categorical and localisation labels for targets of interest, such as colonic diseases [54, 57, 62], accessory instruments [42, 34, 42], polyps [48, 59, 23, 64, 72], endoscopic artefacts [45, 36], and gastrointestinal diseases [56, 43]. In addition, the organisation of competitions has accelerated growth within the colonoscopy community by establishing shared platforms for data collection and model evaluation, significantly advancing research areas such as the detection of intestinal disease [40] and polyp [61, 69, 38].
$\bullet$ Segmentation datasets for colonoscopy research originate from two sources. The first source comprises real data, mainly utilised for single-target segmentation of entities such as polyps [16, 30, 32, 33, 44, 49], Crohn’s disease [53], and accessory tools [50]. Some datasets, like BKAI-Small/Large [46], provide instance-level masks for neoplastic and non-neoplastic polyps. Other datasets come from organised competitions, such as polyp segmentation datasets [19, 20, 21, 22, 61] and a gastrointestinal diseases segmentation dataset [40], or extensions of existing databases, offering pixel-wise masks (e.g., for polyps [38, 55], colorectal disease [62]) or scribble labels (e.g., for polyps [59, 67]). The other source, such as SinGAN-Seg [63], generates synthetic images for polyp segmentation.
$\bullet$ VL datasets remain relatively scarce so far, with two known datasets for this specific purpose. MEDVQA-GI [65] contributes the first dataset with three VL tasks, including visual question answering, visual question generation, and visual location question answering. Kvasir-VQA [74] collects 6,500 question-answer pairs from existing datasets [43, 50], for gastrointestinal diagnostics, such as image captioning, visual question answering.
3.2 Discussion
Based on the 63 revisited datasets, we have several data-centric observations that could inspire more future ideas.
$\bullet$ Data granularity requires improvement to better understand patient conditions and treatment efficacy. (a) More than a quarter of polyp-containing datasets provide fine-grained categorisation, often with inadequate detail. For example, BKAI-Small/-Large [46] provide two instance annotations of neoplastic and non-neoplastic polyps. ColonoscopicDS [26] categorises at the video level into hyperplasic, serrated, and adenoma lesions. SUN-database [54] provides fine-grained labels, documenting measurements of polyp size (height, width) and morphology (pedunculated, sessile, flat), along with their anatomical locations (e.g., rectum, sigmoid colon). Several data-centric areas remain under-explored, such as temporal lesion evolution recording, granularity improvement, graded severity tagging, and instance-level target annotation. (b) Furthermore, label orthogonality, an often overlooked issue, treats labels as isolated entities. Current works seldom discuss potential inter-class correlations, such as the co-occurrence of inflammatory bowel disease with erosion symptoms, Crohn’s disease with fistula complications, or colorectal cancer accompanied by bleeding. Future studies should consider causality [88] and comorbidity [89] to address these correlations effectively.
$\bullet$ Data diversity is crucial to developing fair and reliable models. Three aspects deserve consideration. (a) Datasets for rare colorectal diseases appear to be limited, due to case scarcity and expertise requirements. For example, Crohn’s disease, which affects an estimated 58 to 241 per 100,000 adults in the United States [90], has so far been discussed in three datasets [53, 58, 52]. Such an unbalanced situation leads to data-hungry models performing better in common cases than in rarer or novel ones. Thus, increasing attention to rare gastrointestinal diseases could potentially improve the ability to treat long-tailed [91] or open-vocabulary [92] problems. (b) Multimodal research in colonoscopy appears to be in its early stages, with limited data [65, 74] for analysis. Therefore, collecting varied patient information (e.g., age, gender, eating habits) combined with expert interpretations (e.g., clinical report, medication advice) could be advantageous and ultimately facilitate personalised and side-effect-minimised colonoscopy practices [93].
$\bullet$ Data inconsistency in colonoscopy research is due to two main factors. (a) Expert interpretations vary due to varying experience, expertise, and observed nuances, leading to subjective judgements and labelling uncertainties. For example, ColonoscopicDS [26] provides raw diagnostic labels for each sample from multiple experts and beginners, reflecting their underlying (dis)agreement. In addition, SUN-SEG [62] releases the rejected segmentation masks from their annotation workflow, highlighting the challenges in reaching consistent polyp boundaries. (b) Existing datasets often have study-specific targets, leaving others unlabelled or classified as background. Nerthus [28], for example, focuses on assessing the quality of bowel preparation, but ignores diagnostic findings such as polyps. Furthermore, multiple categories in GastroVision [66] are not mutually exclusive; for example, a case labelled as “accessory tool” could also fall into the category “blood in lumen”. Segmentation data like Kvasir-Instrument [50] annotates only medical instruments, ignoring other targets like polyps, whereas Kvasir-SEG [38] provides polyp label, leaving out others like instruments. These observations prompt future research into learning from partial [94], noisy [95], or missing [96] labels.
TABLE II: Summary of classification models in colonoscopy. Dataset: CU=CU-ColonDB [25], CDS=ColonoscopicDS [26], Private=private data, HK= HyperKvasir [43], KC=Kvasir-Capsule [56]. Backbone: CaffeNet [97], D-121=DenseNet121 [98], R-12/-18/-50/-101=ResNet12/18/50/101 [99], ViT-S16 or ViT-B16 [100], MobV2=MobileNetV2 [101], R50-Att=ResNet50 with attention module [102], C3D [103], Inc-v3=Inceptionv3 [104], I3D [105]. “Customised” means a base network modified for the current task or a model independent of the base network choice. Head: classifier implemented by the fully connected (FC) and support vector machine (SVM) layers, or using the $\ell^{2}$ norm to measure the disparity between the input and output. Arch: the architectures shown in Fig. 3. Sup: learning strategies such as fully supervised (FS), semi-supervised (SS), unsupervised (US), and weakly supervised (WS). For simplicity, the following tables use consistent abbreviations unless specified otherwise.
Model Publication Core design Training dataset Testing dataset Backbone Arch Head Sup URL
Image-based models Zhang et al. [25] JBHI’16 domain transfer learning CU, CDS CU, CDS CaffeNet BF#1 SVM FS -
RIIS-DenseNet [106] MICCAI’18 rotation-invariant, similarity constrained Private Private D-121 SF FC FS -
FSAD-Net [107] MICCAI’20 mutual information maximisation Private Private D-121 BF#2 FC US Link
Gammulle et al. [108] MICCAI’20 relational mapping Kvasir [28], Nerthus [29] Kvasir [28], Nerthus [29] R-50 MF#1 FC FS -
ADGAN [37] ISBI’20 dual adversarial learning Liu et al.[37] Liu et al.[37] Customised BF#2 $\ell^{2}$ US -
Carneiro et al. [109] MedIA’20 model uncertainty & calibration Private Private D-121 SF FC FS -
SSL-WCE [110] MedIA’20 adaptive aggregated attention CAD-CAP [41] CAD-CAP [41] D-121 BF#2 FC SS Link
PolypsAlign [47] MICCAI’21 teacher-student alignment CPC-Paired [47] CPC-Paired [47] R-50 BF#2 FC FS Link
CPC-Trans [111] MICCAI’22 cross-modal representation consistency CPC-Paired [47] CPC-Paired [47] ViT-S16 BF#2 FC FS Link
FFCNet [112] MICCAI’22 frequency domain learning Private Private R-18 SF FC FS Link
DLGNet [113] MedIA’23 Gaussian mixture model Private Private R-18 BF#2 FC FS Link
Yue et al. [114] TIM’23 class imbalance loss Private, HK Private, HK MobV2 SF FC FS Link
DAFON [115] ESWA’24 few-shot open-set learning Kvasir-Capsule [56] Kvasir-Capsule [56] R-12 BF#2 FC FS -
SSL-CPCD [73] TMI’24 composite pretext-class discrimination LIMUC [68] Private, LIMUC [68] R50-Att BF#2 FC FS Link
Video BseNet [116] MICCAI’18 unsupervised depth estimation, LSTM [117] Private Private C3D SF FC FS -
Byrne et al. [118] Gut’19 real-time assessment system Private Private Inc-v3 SF FC FS -
Tamhane et al. [119] MICCAIw’22 vision transformer based Private Private ViT-B16 SF FC FS -
Tian et al. [59] MICCAI’22 multiple instance learning WVAD [59] WVAD [59] I3D SF FC WS Link
Figure 3: Gallery of deep-based architectures. The single-stream framework (SF) features a single input and output with sequential data flow. Multi-stream frameworks predict a single output but involve parallel processing streams, either at the decoding stage (MF#1) or the encoding stage (MF#2). Branched frameworks extend multi-stream framework to produce multiple outputs from either a single input (BF#1) or multiple inputs (BF#2). These side outputs typically receive supervision from additional supervisory signals, such as boundary cues.
4 Revisiting Colonoscopy Models
In this section, we investigate 137 deep learning models for colonoscopic scene perception, sourced from leading conferences or journals published since 2015. First, three widely recognised topics are described, including 18 classification models (§4.1), 24 detection models (§4.2), and 86 segmentation models (§4.3). Their architectural designs are classified into five subtypes as presented in Fig. 3. Lastly, we discuss nine VL-related models in §4.4.
4.1 Classification models
$\bullet$ Input phase. Tab. II lists the training and testing data used for each deep model. We note that many classification models [106, 109, 112, 114, 113, 116, 118] in colonoscopy use in-house data for model development, partly resulting in the absence of domain-recognised benchmarks. This issue stems from the different categorical goals pursued by individual studies, such as identifying polyps from white-light and narrow-band imaging pairs [47, 111], images [106], or videos [59], evaluating polyp size [116], and recognising colonic diseases [112] or landmarks [119].
$\bullet$ Processing phase. We discuss data flow management strategies based on two attributes. (a) Backbone: Early models [108, 106] typically employ well-trained convolutional backbones (e.g., ResNet [99], DenseNet [98]) from ImageNet [3], while recent studies explore alternatives such as using a vision transformer in [111] and a lightweight network in [114]. Another strategy is SSL-CPCD [73], which involves pre-training a model to generate domain-specific representations, followed by its generalisation on various downstream perception tasks. (b) Architecture: Classification models involve various designs as in Fig. 3. At first, a basic idea is of using a single-stream framework (SF) that sequentially processes visual features based on confidence calibrated [109] or 3D convolutional [116, 59] networks. Second, Gammulle et al.[108] proposed a dual decoding flow approach for hierarchical feature encoding, typified as MF#1. Third, to ensure reliable predictions, branched frameworks are adopted for multi-objective learning, such as the integration of parallel feature flows [25], interclass Gaussian loss [113], and global-to-local consistency [111].
TABLE III: Summary of detection models in colonoscopy. Dataset: C6=CVC-ClinicDB [21], ETIS=ETIS-Larib [20], ASEI [34], C3=CVC-ColonDB [19], KUMC [57], LDPV=LDPolyVideo [48], SUN=SUN-database [54], PL=PICCOLO [39], KID=KID1&2 [31], CDS [26], KSe=Kvasir-Sessile [55], ASU =ASU-Mayo [22], CDB=CVC-ClinicVideoDB [27], ES=EndoSceneStill [30], CU [25], ACP=ACP-ColonDB530 [42]. Backbone: R-34/-50/-101 [99], CDN-53=CSPDarkNet53 [120], DN-53=DarkNet-53 [121], EffDet-D0=EfficientDet-D0 [122], AlexNet [123], V-16=VGG16 [124], R-50v2= ResNet50V2 [125]. WF: one-stage (OS) or two-stage (TS) workflows. NMS: non-maximum suppression. EC: edge-sensitive cues.
Model Publication Core design Training dataset Testing dataset Backbone WF Arch NMS EC Sup URL
Image-based models Yang et al. [126] TIM’20 parallel detection & segmentation Private, C6, ETIS Private, C6, ETIS R-50 TS BF#1 FS -
ConsolidatedPolypDA [127] MedIA’21 Gaussian Fourier domain adaptation C6 ETIS, ASEI R-101 TS BF#2 US Link
MDeNetplus [128] MedIA’21 2D Gaussian shapes prediction C6 C3, ETIS R-34 OS MF#1 FS -
FedInI [129] MICCAI’22 federated learning, structural causal model KUMC KUMC R-101 TS BF#2 FS Link
Pacal et al. [130] CIBM’22 improved YOLOv3 [121]/v4 [120] SUN, PL SUN, PL, ETIS CDN-53/DN-53 OS BF#1 FS Link
SMPT++ [131] TMI’22 source-free domain adaptation Private, C6, ETIS, ASEI, KID Private, C6, ETIS, ASEI, KID R-101 OS BF#1 US Link
FRCNN-AA-CIF [132] CIBM’23 attention module & context information fusion Private Private R-101 TS BF#1 FS -
Haugland et al. [133] MI’23 modality translation Private, PL, CDS PL, KUMC EffDet-D0 OS BF#2 FS -
SCAN++ [134] TMM’23 enhanced semantic conditioned adaptation C6, ASEI C6, ASEI R-101 OS BF#2 FS, US Link
TFCNet [135] CIBM’24 fine-grained feature compensation C6, KUMC, LDPV C6, KUMC, LDPV, KSe CDN-53 OS BF#1 FS -
DUT [136] TNNLS’24 decoupled unbiased teacher C6, ASEI, Private ASEI, Private R-101 OS BF#2 US Link
Video-based models Tajbakhsh et al. [137] IPMI’15 patch descriptor & edge classification Private Private AlexNet TS BF#1 FS -
Tajbakhsh et al. [22] TMI’15 extension on [137] C3 C3, ASU AlexNet TS BF#1 FS -
Yu et al. [138] JBHI’16 online and offline integration ASU ASU Customised OS MF#2 FS -
Mo et al. [139] ICPR’18 building upon Faster R-CNN [140] CDB C6, C3, CDB, ES V-16 TS BF#1 FS -
Qadir et al. [141] JBHI’19 temporal dependency ASU, C6 ASU, CDB V-16 TS BF#1 FS -
AIPDT [142] MICCAI’20 parallel detection and tracking Private, CDB CDB DN-53, AlexNet OS BF#2 FS -
AI-doscopist [42] NPJDM’20 spatial-temporal fusion C6, ETIS, C3, ASU, CU, ACP C6, ETIS, C3, ASU, CU, ACP R-50 OS BF#2 FS -
STFT [143] MICCAI’21 spatial-temporal feature transformation ASU, CDB ASU, CDB R-50 OS BF#2 FS Link
Yu et al. [144] AIM’22 instance tracking head (plug-and-play) Private, C6, CDB Private, CDB, ETIS V-16 OS BF#2 FS -
EMSEN [145] TII’22 explainable multitask Shapley explanation CDS CDS Customised OS BF#2 FS -
YONA [146] MICCAI’23 feature alignment & contrastive learning SUN, CDB, LDPV SUN, CDB, LDPV R-50 TS BF#2 FS Link
Intrator et al. [147] MICCAI’23 self-supervised polyp re-identification Private Private R-50v2 OS MF#2 US -
V2I-DETR [148] arXiv’24 video-to-image knowledge distillation SUN SUN R-50 OS BF#2 FS -
$\bullet$ Output phase. (a) Prediction head: An early model [25] applies two SVM layers to classify polyps into three categories. Modern methods usually adopt a fully connected layer as the final classifier due to its simplicity and flexibility. A special case is ADGAN [37], a generative adversarial network that computes the $\ell^{2}$ norm differential between input and output to identify anomalous images. (b) Supervision strategy: Most methods use fully-supervised learning with pre-annotated categories, but some explore data-efficient strategies, including semi-supervised [110], weakly-supervised [59], and unsupervised [37, 107] learning.
$\bullet$ Remarks. We observe three aspects of the above classification models. (a) Novel visual backbones, like the state space model [149], remain underexplored. Furthermore, reformulating the classification paradigm within VL models, e.g., CLIP [150], can yield unexpected results. (b) Benchmarks for multicategory classification remain underexplored. The Kvasir series [28, 43, 56] offers valuable sources for further study. We will explore these public data on their potential synergy in §5.1. (c) Several new task settings have emerged in colonoscopy. For example, Tian et al.[59] recognise abnormal frames from colonoscopy videos from an out-of-distribution view. DAFON [115] solve an open-set classification problem within a few-shot framework.
4.2 Detection models
$\bullet$ Input phase. Detection models classify targets and locate them using boxes, assisting surgical intervention and planning. The goals of interest are diverse, such as identifying the polyp(s) in images [126, 127, 128, 130, 131, 132, 133, 134, 136]/videos [137, 22, 138, 139, 141, 142, 143, 144, 146, 145, 147, 148], or locating multiple findings [42] like bleeding, polyps, and accessory tools.
$\bullet$ Processing phase. This has three key configurations for the analysis. (a) Backbone: There are two general strategies for network initialisation. The first group [126, 127, 128, 131, 132, 134, 136, 42, 143, 146, 147, 148] leverages the ResNet series [99] pre-trained on the ImageNet [3] dataset. The second group relies on well-trained object detectors, such as [130, 142] using DarkNet series [121, 120] and [133] employing EfficientDet-D0 [122]. (b) Workflow: Detection models are often built on general-purpose architectures. In the “WF” column of Tab. III, we categorise the models according to their processing stages. The two-stage workflow decouples detection into the region proposal and classification phases, like models [127, 132, 139, 141] based on Faster R-CNN [140]. One-stage models prioritise speed and simplicity, operating in a single forward. For example, some studies [130, 142, 42] are based on the YOLO framework [121, 120], and Yu et al. [144] uses the SSD framework [151]. (c) Architecture: Detection models predict target categories and spatial coordinates, typically implemented in branched frameworks (BF#1/BF#2) as shown in Fig. 3. Two special cases [128, 138] adapt the design of multistream framework to first pop out pixel-wise attention regions, then convert them to bounding boxes.
TABLE IV: Summary of segmentation models in colonoscopy. Dataset: C6 [21], ES [30], KS=Kvasir-SEG [38], C3 [19], ETIS [20], HK [43], ASU [22], CDB [27], BKAI=BKAI-Small [46], KSe [55], GI=GIANA [152], SUN-S [62], PG=PolypGen [69], K-I=Kvasir-Instrument [50]. Backbone: ResUNet [153], R-18/-34/-50/-101 [99], R-50v2 [125], R2-50=Res2Net-50 [154], V-16 [124], DeiT [155], Eff-B4= EfficientNet-B4 [156], DLV3+=DeepLab V3+ [157], PB2/3/5=PVTv2-B2/-B3/-B5 [158], CvT [159], MiT-B2 [160], CMLP=CycleMLP [161], P-DETR=Point DETR [162], D-121=DenseNet121 [98], MN= MSCAN [163], Swin-T [164], SAM [165], SAM2 [166], ViT-B16 [100], DLV2=DeepLabV2 [167], HR-W48=HRNet-W48 [168], CN-T=ConvNeXt-Tiny [169], SFB3=SegFormer-B3 [160], M2Former=Mask2Former [170]. Edge-sensitive analysis by explicitly (EX) using edge map as supervision or input and implicitly exploring edge-aware representation (IM#1) or edge-aware uncertainty (IM#2).
Model Publication Core design Training dataset Testing dataset Backbone Arch Edge Sup URL
Image-based models Yuan et al. [171] JBHI’17 weak bottom-up & strong top-down saliency Private C6 Customised BF#1 - US -
SFA [172] MICCAI’19 area & boundary constraints ES ES Customised BF#1 EX FS Link
ResUNet++ [173] ISM’19 enhanced deep residual UNet [153] C6, KS C6, KS ResUNet MF#1 - FS Link
ACSNet [174] MICCAI’20 adaptive context selection ES, KS ES, KS R-34 BF#1 - FS Link
PraNet [175] MICCAI’20 reverse attention, parallel partial decoder C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 EX, IM#1 FS Link
UI-CNN [176] MedIA’20 Monte Carlo guided back-propagation ES ES V-16 MF#1 IM#2 FS Link
ThresholdNet [44] TMI’20 confidence-guided manifold mixup ES ES, WCE-Polyp [44] R-101 BF#2 EX, IM#1 FS Link
SCR-Net [177] AAAI’21 semantic calibration & refinement KS KS Customised MF#1 - FS Link
BI-GCN [178] BMVC’21 boundary-aware graph convolution C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 EX FS Link
FDSemi [179] ICCV’21 collaborative & adversarial learning C6, KS C6, KS Customised BF#2 IM#1 SS Link
CCBANet [180] MICCAI’21 cascading context & balancing attention C6, ES, KS C6, ES, KS R-34 BF#1 IM#1 FS Link
CCD [181] MICCAI’21 constrained contrastive distribution learning HK, Liu et al.[37] HK, Liu et al.[37] R-18 BF#2 - US Link
HRENet [182] MICCAI’21 hard region enhancement C6, KS C6, KS, C3 R-34 BF#1 EX FS Link
LOD-Net [183] MICCAI’21 learnable oriented derivatives C6, KS C6, ES, KS, C3, ETIS R-50 MF#2 EX FS Link
MSNet [184] MICCAI’21 multiscale subtraction network C6, KS C6, ES, KS, C3, ETIS R2-50 MF#1 EX, IM#1 FS Link
SANet [185] MICCAI’21 colour exchange & probability correction C6, KS C6, ES, KS, C3, ETIS R2-50 MF#2 - FS Link
Transfuse [186] MICCAI’21 fusing transformers and CNNs C6, KS C6, ES, KS, C3, ETIS R-50v2, ViT-B16 BF#2 EX FS Link
EndoUDA [49] MICCAI’21 domain adaptation, variational autoencoder training Celik et al.[49] Celik et al.[49] Eff-B4 BF#2 EX FS, US Link
UACANet [187] MM’21 uncertainty augmented context attention C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 EX, IM#2 FS Link
Jha et al. [55] JBHI’21 ResUNet++ [173] with conditional random field & test-time augmentation C6, C3, ETIS, KS, ASU, CDB C6, C3, ETIS, KS, ASU, CDB ResUNet MF#1 - FS Link
MPA-DA [188] JBHI’21 mutual-prototype adaptation network ES, KS ETIS R-101 BF#2 IM#2 FS, US Link
DW-HieraSeg [189] MedIA’21 hierarchical segmentation, dynamic-weighting ES ES DLV3+ BF#1 - FS Link
ICGNet [190] IJCAI’22 context-based reverse-contour guidance ES, KS ES, C3, KS R-34 BF#1 EX FS -
BoxPolyp [191] MICCAI’22 segmentation with extra box labels C6, KS C6, ES, KS, C3, ETIS R2-50, PB2 BF#1 - WS Link
LDNet [192] MICCAI’22 dynamic kernel generation & update Private, C6, KS Private, C6, KS, C3, ETIS R2-50 BF#1 - FS Link
PPFormer [193] MICCAI’22 polyp-guided self-attention, local-to-global method C6, KS C6, ES, KS, C3, ETIS V-16, CvT BF#1 EX FS -
SSFormer [194] MICCAI’22 stepwise local & global feature aggregation C6, KS C6, C3, ETIS, KS MiT-B2 MF#1 - FS Link
TRFR-Net [195] MICCAI’22 task-relevant feature replenishment C3, ETIS, KS C3, ETIS, KS R-34 BF#2 - FS, US Link
AFP-Mask [196] JBHI’22 anchor-free instance segmentation Private, GI Private, C6, ETIS Customised BF#1 - FS -
BCNet [197] JBHI’22 cross-layer integration, bilateral boundary extraction KS C6, ES, KS R2-50 BF#1 EX FS -
BSCA-Net [198] PR’22 bit slice context attention C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 EX FS Link
Polyp-Mixer [199] TCSVT’22 context-aware MLP-based network C6, KS C6, KS, C3, ETIS CMLP BF#1 - FS Link
ACL-Net [200] AAAI’23 affinity contrastive learning C6, KS C6, ES, KS, C3, ETIS R-50 BF#2 - SS Link
WS-DefSegNet [67] CVPRw’23 deformable transformer, sparse foreground loss W-Polyp [67] C6, KS, C3, ETIS, ES R2-50 BF#1 - WS, SS Link
WeakPolyp [201] MICCAI’23 mask-to-box transformation, scale consistency SUN-S, Private SUN-S, Private PB2 BF#2 - WS Link
PETNet [202] MICCAI’23 Gaussian-probabilistic guided semantic fusion C6, KS C6, ES, KS, C3, ETIS PB2 BF#1 - FS Link
S2ME [203] MICCAI’23 spatial-spectral mutual teaching, ensemble learning SUN-S C6, KS, SUN-S, PG Customised BF#1 - WS Link
Su et al. [204] MICCAI’23 feature propagation & aggregation C6, KS C6, ES, KS, C3, ETIS PB2 BF#1 EX FS -
Polyp-PVT [205] AIR’23 Improved pyramid vision transformer C6, KS C6, ES, KS, C3, ETIS PB2 BF#1 EX FS Link
RPANet [206] IPMI’23 coarse-to-fine self-supervision Private C6, ETIS, KS R-101 BF#2 - FS, US -
FEGNet [207] JBHI’23 feedback enhancement gate network C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 EX FS -
BS-Loss [208] JBHI’23 boundary-sensitive loss with location constraint K-I K-I Customised MF#1 EX FS Link
Point SEGTR [209] MedIA’23 multi-point and symmetric consistency C6, ETIS C6, ETIS P-DETR BF#2 - FS, WS, SS -
DGNet [210] MIR’23 deep gradient learning C6, KS C3, ETIS Eff-B4 BF#1 EX FS Link
CFA-Net [211] PR’23 cross-level feature fusion, boundary aggregation C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 EX FS Link
ColnNet [212] TMI’23 statistical attention, anomaly boundary approximation C6, KS C6, ES, KS, C3, ETIS D-121 BF#1 EX FS -
FANet [213] TNNLS’23 feedback attention, run-length encoding C6, KS C6, KS Customised MF#2 - FS Link
MCANet [214] arXiv’23 multi-scale cross-axis attention C6, KS C6, ES, KS, C3, ETIS MN MF#1 - FS Link
Polyper [215] AAAI’24 boundary sensitive attention C6, KS C6, KS, C3, ETIS, ES Swin-T MF#1 IM#1 FS Link
EMCAD [216] CVPR’24 efficient multi-scale convolutional attention decoding C6, C3, ETIS, KS, BKAI C6, C3, ETIS, KS, BKAI PB2 BF#1 - FS Link
Schön et al.[217] CVPR’24 SAM [165], test-time adaptation K-I, CDB, KS K-I, CDB, KS SAM BF#2 EX WS -
MH-pFLID [218] ICML’24 federated learning, injection-distillation paradigm Private Private Customised BF#2 - FS Link
ASPS [219] MICCAI’24 SAM [165], uncertainty-guided regularisation C6, KS C6, ES, KS, C3, ETIS SAM, MN BF#2 IM#2 FS Link
Polyp-Mamba [220] MICCAI’24 vision state space model, semantic relationship mining C6, KS C6, ES, KS, C3, ETIS VMamba [221] MF#1 EX FS -
QueryNet [222] MICCAI’24 unified framework of segmentation & detection C6, KS C6, ES, KS, C3, ETIS M2Former MF#2 - FS Link
LSSNet [223] MICCAI’24 local & shallow feature supplementation C6, KS C6, ES, KS, C3, ETIS PB2 BF#1 EX FS Link
BSBP-RWKV [224] MM’24 Perona-Malik diffusion, RWKV [225] KS KS Customised BF#1 EX FS -
CFATransUnet [226] CIBM’24 channel-wise cross fusion attention and transformer C6, KS C6, KS PB3 BF#1 - FS Link
PolypMixNet [227] CIBM’24 consistency regularisation, soft pseudo labeling C3, C6, KS, ETIS C3, C6, KS, ETIS R-34 BF#2 - SS Link
RGIAug [228] JBHI’24 randomised global illumination augmentation C3, C6, ETIS, KS C3, C6, ETIS, KS Customised BF#2 - FS Link
EMTS-Net [229] JBHI’24 multi-task synergetic network C6, KS C6, ES, KS, C3, ETIS R2-50 BF#1 - FS -
MSDE-Net [230] JBHI’24 multi-scale dual-encoding network K-I K-I R-34, DeiT MF#2 - FS -
Polyp-OOD [231] MIR’24 out-of-distribution modelling, latent standardisation SUN-S SUN-S, C6, C3, ETIS, KS ViT-B16 MF#2 - US Link
MedSAM [232] NComs’24 SAM [165], cross-organ/modality tuning Hybrid datasets Hybrid datasets SAM MF#2 - WS Link
FoBS [233] TCSVT’24 multi-level boundary-enhanced framework KS, ES KS, ES, ETIS, C3 DLV3+ BF#2 EX, IM#1 FS Link
DCL-PS [234] TMI’24 domain-interactive contrastive learning, self-training ETIS, HK, ES, KS ES, KS DLV2 BF#2 - FS, US Link
Gao et al.[235] TMI’24 in-context learning, dual similarity checkup C3 C3 SAM BF#2 - WS
U-KAN [236] arXiv’24 U-shaped Kolmogorov-Arnold network [237] C6 C6 Customised MF#1 - FS Link
SliceMamba [238] arXiv’24 vision state space model, bidirectional slice scan C6, KS C6, KS Customised MF#1 - FS -
ProMamba [239] arXiv’24 vision state space model, promptable segmentation C6, KS C6, KS, C3, ETIS, ES, BKAI Vim [149] MF#2 - WS -
SAM2-UNet [240] arXiv’24 SAM2 [166], adapter-based tuning C6, KS C6, ES, KS, C3, ETIS SAM2 BF#1 EX FS Link
Video-based models Puyal et al. [241] MICCAI’20 temporal correlation via hybrid 2D/3D CNNs Private, KS Private R-101 MF#2 - FS -
PNS-Net [242] MICCAI’21 normalised self-attention, progressive learning C6, C3, ASU, KS C6, C3 R2-50 BF#2 - FS Link
SSTAN  [243] MICCAI’22 spatial-temporal attention LDPolyVideo [48] LDPolyVideo [48] ResUNet BF#2 - SS Link
TCCNet [244] IJCAI’22 temporal consistency, context-free loss C6, C3 C6, C3, ETIS R2-50 BF#2 EX, IM#1 SS Link
Puyal et al. [245] MedIA’22 extend [241] with optimal setups Private, KS Private, SUN R-101 BF#2 - FS -
PNS+ [62] MIR’22 extend [242] with global-to-local learning SUN-S SUN-S R2-50 BF#2 - FS Link
EUVPS [246] AAAI’24 cross-scale region linking, cross-wise scale alignment SUN-S, C6 SUN-S, C6 HR-W48 BF#2 - FS Link
LGRNet [247] MICCAI’24 cyclic neighbourhood propagate, Hilbert selective scan C6, C3, SUN-S C6, C3, SUN-S R2-50 BF#2 - FS Link
SALI [248] MICCAI’24 short-term alignment, long-term interaction module SUN-S SUN-S PB5 BF#2 EX FS Link
Diff-VPS [249] MICCAI’24 diffusion model, adversarial temporal reasoning SUN-S SUN-S SFB3 BF#2 - FS, US Link
FlowICBNet [250] CIBM’24 iterative feedback units, frame filtering & selecting SUN-S SUN-S R2-50 BF#2 - FS Link
Drag&Drop [251] MIR’24 weakly-supervised temporal annotator SUN-S SUN-S - BF#2 - WS Link
SSTFB [252] arXiv’24 self-supervised encoder, sub-branching mechanism SUN-S SUN-S, CDB R2-50 BF#2 - US, FS -
Vivim [253] arXiv’24 video state space model, spatio-temporal selective scan KS, ASU, C3, C6 C3, C6 Customised BF#2 EX FS Link
MAST [254] arXiv’24 Siamese transformer, mixture attention module SUN-S SUN-S PB2 BF#2 - FS Link
$\bullet$ Output phase. (a) Post-processing techniques are employed to eliminate duplicate predictions and select the most relevant targets, with non-maximum suppression (NMS) being a widely-adopted method [130, 131, 132, 133, 139, 141, 42, 144, 146]. (b) Auxiliary information can improve prediction reliability, such as edge cues suggested to provide geometric patterns for object detection in [128, 137, 22]. (c) Supervision strategy is currently dominated by fully supervised learning, such as using region-level labels [133, 130, 126, 132, 139, 141, 148, 143, 145, 142, 144] and pixel-wise [138, 128, 137, 22] labels, and introducing the box-assisted contrastive learning [146]. Other models [127, 131, 134, 136] explore unsupervised domain adaptation techniques to detect polyps across colonoscopy devices.
$\bullet$ Remarks. We emphasise a few observations for the above review. (a) Most models focus on detecting polyp(s), while other colonoscopic findings receive less attention. We encourage exploring public multitarget [34] or multicentre [78] data. (b) Learning strategies are underexplored. General-purpose detection models like using weak supervision [255] provide valuable references, being potentially more feasible and cost-effective since they require less detailed annotations from medical experts. (c) Beyond well-established convolution-based detection frameworks, recent methods like transformer-based DETR [256] and diffusion-based DiffusionDet [257], open exciting opportunities for this field. Moreover, exploring cross-task synergy is promising, as three video-based models [142, 144, 147] demonstrated effectiveness in unifying polyp detection and tracking frameworks. (d) Although some datasets such as the SUN-database [54] ($>$158K samples) and LDPolyVideo [48] ($>$900K samples) are relatively large, this field still lacks a standardised evaluation benchmark.
TABLE V: Comparison of image polyp segmentation models. They are evaluated using the mean scores (%) of structure measure ($\mathcal{S}$ [258]) and Dice coefficient ($\mathcal{D}$) on two test sets, with boxplots illustrating the distribution of their consistency and variability across test cases.
Kvasir-SEG (100 test images) [38] CVC-ClinicDB (62 test images) [21]
[…]
Polyp-PVT [205] 92.51 [#1] [Uncaptioned image] 91.74 [#2] [Uncaptioned image] 95.00 [#2] [Uncaptioned image] 93.68 [#1] [Uncaptioned image]
CFA-Net [211] 92.40 [#2] 91.47 [#4] 95.07 [#1] 93.25 [#2]
MSNet [184] 92.31 [#3] 90.23 [#7] 94.68 [#3] 91.48 [#6]
BoxPolyp [191] 92.30 [#4] 91.84 [#1] 93.70 [#6] 91.81 [#4]
SSFormer [194] 92.21 [#5] 91.71 [#3] 92.87 [#9] 90.60 [#7]
UACANet [187] 91.67 [#6] 91.21 [#5] 94.30 [#4] 92.63 [#3]
PraNet [175] 91.50 [#7] 89.82 [#8] 93.68 [#7] 89.90 [#9]
SANet [185] 91.45 [#8] 90.41 [#6] 93.98 [#5] 91.57 [#5]
DGNet [210] 90.98 [#9] 89.72 [#9] 93.39 [#8] 90.44 [#8]
MCANet [214] 90.25 [#10] 89.55 [#10] 91.79 [#10] 89.70 [#10]
Polyper [215] 90.08 [#11] 89.12 [#11] 91.29 [#11] 88.63 [#11]
UNet++ [259] 86.21 [#12] 82.08 [#12] 87.33 [#13] 79.42 [#13]
UNet [260] 85.76 [#13] 81.83 [#13] 89.00 [#12] 82.25 [#12]
SFA [172] 78.14 [#14] 72.31 [#14] 79.33 [#14] 70.06 [#14]
4.3 Segmentation models
Compared to the above two topics, the segmentation research appears to be well established. This is evidenced by the extensive amount of research documented in Tab. IV.
$\bullet$ Input phase. Most segmentation models focus on a single target (i.e., polyp), typically adopting as binary segmentation paradigm. These models usually follow the well-established testbed of PraNet [175] for their development and comparison. An exception case, AFP-Mask [196], provides an anchor-free framework for segmenting polyp instances. Recent works [208, 230, 217] have also focused on segmenting surgical tools during procedures.
$\bullet$ Processing phase. (a) Backbone: The visual encoders for the segmentation task have been extensively explored. A common option is to use a general backbone pre-trained on ImageNet [3], such as using CNN [175, 214], vision Transformer [205, 215], hybrid CNN-Transformer network [223, 193, 186], multilayer perceptron [199], state space model [220, 239, 253, 238], receptance weighted key value (RWKV) [224], and Kolmogorov-Arnold network [224]. An alternative is to use well-trained perception models such as Point DETR [162] used in [209], DeepLab series [157, 167] applied in [189, 233, 234]. Recently, there has been a shift towards promptable architectures. The first way is to exploit the foundation models, for example, by fine-tuning the segment anything modal (SAM) [165] with location prompts [232], adapting SAM during the test time [217], exploiting the hybrid CNN-Transformer network [219], or incorporating trainable adapter layers into the SAM2’s encoder [240]. Another way aims to adapt the model to unseen scenarios through in-context learning [235]. (b) Architecture: The community favours the encoder-decoder design for its superior ability to perceive hierarchical features. Current models usually opt for multistream or branched frameworks, as in Fig. 3. Various modifications in this area have been explored, such as incorporating residual connected flows [173, 55], probing cross-task synergy [222], providing additional edge cues [210], using the model-ensembling strategy [186, 193, 219], calculating latent statistics [231], exploring spatio-temporal relationships through 3D convolutions [241, 245] or self-attention modules [242, 62, 248], and approaching with the teacher-student paradigm [67, 209]. (c) Edge-sensitive analysis: Geometric patterns are beneficial in enhancing the model’s capability to differentiate foreground objects from the background. The current techniques are in two main ways. The first involves the explicit use of edge maps derived from image gradients, either for direct supervision [172, 182, 210] or as an auxiliary input [49]. Moreover, some methods emphasise edge-aware calculation within their loss functions, such as boundary weighted [193, 240] and customised [183, 208]). Second, edge information can be implicitly integrated by embedding edge-aware representations (e.g., reverse attention [175], morphology operator [215], subtraction operator [44]), or by quantifying edge-aware uncertainties [176, 187]. Moreover, some methods adopt a hybrid strategy, e.g., both the subtraction operator and the edge-aware loss are used in MSNet [184].
$\bullet$ Output phase. Most models are trained in a fully supervised way, always incorporating deep supervision at various decoding stages, as seen in [175, 205]. Recent models have extended beyond with data-efficient approaches, for example, weakly supervised mask-to-box transformation [201] and unsupervised techniques such as contrastive learning [181], out-of-distribution modelling [231], and pseudo-label supervision [243, 191]. Hybrid supervised strategies are also present in which models [49, 188, 195, 206] undergo fully supervised training in the source domain, and are then adapted to the target domain in an unsupervised way. Point SEGTR [209] exploits three types of supervision to enhance the model. Additionally, some teacher-student networks receive hybrid supervision signals; for example, Ren et al.[67] employs a weakly supervised approach for the teacher while the student undergoes semi-supervised training.
$\bullet$ Remarks. To reflect current field progress, we analyse 14 open-source image segmentation models on two popular test datasets, as shown in Tab. V. All models are trained on the same dataset from Fan et al.’s benchmark [175]. First, we observe that current learning strategies are underexplored, as evidenced by BoxPolyp, a weakly supervised model obtaining the best $\mathcal{D}$ score (91.84%) on Kvasir-SEG. In addition, some models achieve better performance, yet exhibit wider interquartile ranges in boxplots, indicating their variability in predictions. For example, Polyp-PVT, which ranks highest in the $\mathcal{S}$ score on Kvasir-SEG, shows a wider interquartile range than other models like SSFormer. From these results, we suggest several promising opportunities for future study. (a) The current gold benchmark [175] comprises less than 1.5K samples and is focused on one category (polyp). In general, scaling up both the data size and diversity could be a natural way to improve robustness and generalisability. This demand is driving innovations, such as developing a semi-auto image annotator [166, 165] to reduce expert labour and synthesising high-fidelity content via diffusion [261] and autoregressive [262] techniques. (b) Moreover, infinite data scaling is not sustainable. Developing data-efficient strategies [94, 95, 96] that require fewer or weaker labels is more cost-effective for average users in the community. (c) Finally, providing procedural support to physicians is essential, including anomaly detection, navigation planning, risk assessment, and intervention advice. We can adopt innovations from similar topics [263].
[…]
Figure 4: Details of the established ColonINST. (a) Three sequential steps to create the instruction tuning dataset for multimodal research. (b) Numbers of colonoscopy images designated for training, validation, and testing purposes. (c) Data taxonomy of three-level categories. (d) A word cloud of the category distribution by name size. (e) Caption generation pipeline using the VL prompting mode of GPT-4V [4]. (f) Numbers of human-machine dialogues created for four downstream tasks.
4.4 VL models
Compared to the above three topics, multimodal research has relatively fewer references. Most existing methods are discriminative models that aim to learn decision boundaries between multimodal inputs. Some studies demonstrate the model’s effectiveness in referring segmentation tasks, for example, by incorporating textual attention of lesion attributes (e.g., size and number of polyps) into a U-shaped model [264], a diffusion model [265], or a hybrid network [266]. Other studies [267, 268] have developd prompt engineering pipelines based on well-trained GLIP [269] for polyp detection. Moreover, the SAM is capable of operating in a VL setting, obtaining location prompts from either the image-text activation map [270] or a zero-shot grounding detector [271, 272]. In the MEDVQA-GI competition [65], most solutions are discriminative-based, approaching VL tasks as a classification mapping problem, where predefined labels are assigned to image-text pairs. An alternative is a generative-based solution [273] that adapts a pretrained BLIP-2 model [274] to generate predictions.
$\bullet$ Remarks. Two possible reasons might explain the lag in VL research for colonoscopy. (a) Data-centric issue. The lack of well-structured and high-quality image-text pairs hinders progress. Future insights can be learnt from existing ideas. First, crawling unlabelled image-text data from social media [275] and scientific literature [276] can be used to build domain-specific foundation models. Second, language models such as GPT-4V [4] can generate diverse professional descriptions, offering an economical and scalable way to expand the knowledge space of the data. (b) Model-centric issue. Current VL techniques in colonoscopy have not kept pace, even with recent developments in multimodal language models (MLMs) [277, 8] for general domains. These techniques opt for a decoder-only strategy that unifies multiple tasks (e.g., detection, captioning) into a unified autoregressive framework (i.e., next-token prediction). These models are flexibly capable of processing input and output texts of varying lengths, without requiring additional task-specific prediction heads for different tasks.
5 Stepping into the Multimodal Land
TABLE VI: Details of instruction tuning dataset ColonINST. For each task, we provide five templates for human instructions, the data sources used to organise human-machine dialogues, and an example of a human-machine conversation.
Task Instruction templates Data source Human-machine dialogue sample
CLS 1. Categorize the object. 2. Determine the object’s category. 3. Identify the category of the object. 4. Classify the object’s category. 5. Assign the object to its corresponding category. 19 sources $\rightarrow$ SUN-database [54], PolypGen [69], CVC-ClinicDB [21], ETIS [20], KUMC [57], Kvasir [28], PSNBI2K [16], CVC-ColonDB [19], EDD2020 [40], Kvasir-Capsule [56], CP-CHILD [51], BKAI-Small [46], PICCOLO [39], WCE-CCDD[60], CPC-Paired [47], HyperKvasir [43], Nerthus [29], GastroVision [66], Kvasi-Instrument [50] [Uncaptioned image]
[…]
CAP 1. Describe what you see in the image. 2. Interpret what the image shows. 3. Detail the visual elements in the image. 4. Explain the image’s visuals thoroughly. 5. Offer a thorough explanation of the image. 19 sources $\rightarrow$ SUN-database [54], PolypGen [69], CVC-ClinicDB [21], ETIS [20], KUMC [57], Kvasir [28], PSNBI2K [16], CVC-ColonDB [19], EDD2020 [40], Kvasir-Capsule [56], CP-CHILD [51], BKAI-Small [46], PICCOLO [39], WCE-CCDD[60], CPC-Paired [47], HyperKvasir [43], Nerthus [29], GastroVision [66], Kvasi-Instrument [50] [Uncaptioned image]
Recently, MLMs have demonstrated significant promise in leveraging language models to process multimodal signals, especially in “perceiving and interpreting” visual signals. Instruction tuning [278] is key to instructing the MLMs to execute domain-specific tasks aligned with user preferences. This section introduces three initiatives to advance multimodal research: how we create a large-scale instruction tuning dataset, ColonINST (§5.1) and how we train a colonoscopy-specific MLM, ColonGPT (§5.2). Finally, as in §5.3, we contribute a multimodal benchmark for conversational colonoscopy tasks, conduct diagnostic studies on ColonGPT, and share lessons from empirical observations.
5.1 Established instruction tuning dataset: ColonINST
$\bullet$ Overview. Fig. (4-a) depicts our semi-automated workflow to create instruction tuning data in three steps. We begin by assembling colonoscopy images from public datasets. Following this, we use a category-specific prompt to interact with a multimodal AI chatbot, GPT-4V [4], yielding descriptive medical captions for these positive cases within the assembled data. Lastly, we reorganise the instruction tuning pairs derived from the data afore-prepared, enabling the model to perform four different colonoscopy tasks in an interactive, dialogue-based manner.
$\bullet$ Data collection. As shown in Fig. (4-b), ColonINST contains 303,001 colonoscopy images, including 128,620 positive and 174,381 negative cases collected from 19 different data sources. To ensure data integrity and avoid label leakage, we establish a series of management rules to divide each dataset. For datasets with predefined divisions, such as KUMC [57], PICCOLO [39], WCE-CCDD [60], BKAI-Small [46], CP-CHILD [51], Kvasir-Instrument [50], and PS-NBI2K [16], we follow their original division rules. When such predefined rules are not available, we conform to widely recognised benchmarks, such as CVC-ClinicDB [21], CVC-ColonDB [19], ETIS-Larib [20], and the polyp category in Kvasir [28] according to the benchmark by Fan et al.[175], as well as positive samples in SUN-database [54] following the benchmark by Ji et al.[62]. For the remaining datasets (HyperKvasir [43], Kvasir-Capsule [56], CPC-Paired [47], Nerthus [29], GastroVision [66], EDD2020 [40], PolypGen [69], negative samples in SUN-database [54], remaining categories in Kvasir [28]), we allocate them proportionally as 60%/10%/30% for training/validation/testing purposes. With the above management rules, our final image division for is roughly 59.73%/8.77%/31.61%. As depicted in Fig. (4-c), all images are classified into a three-level structure, including 4 root/13 parent/62 child categories. In detail, the root level contains three positive categories: the pathological findings of various colonic diseases (110,970 cases); the anatomical structure related to the colon (5,511 cases); and therapeutic interventions related to colonoscopy (12,139 cases), covering the pre-operative, intra-operative, and post-operative stages. Targets outside our interest (e.g., stomach, oesophagus, normal Z-line, and gastric polyp not occurred during a colonoscopy) or lack of objects (e.g., normal mucosa, colon background) are classified under the negative root category (174,381 cases). We intentionally keep these negative samples, as they may be valuable for future gastrointestinal research. Fig. (4-d) presents a word cloud distribution for all categories names within ColonINST.
$\bullet$ Caption generation. The behavioural study [279] suggests that language affects both higher-level (recognition) and lower-level (discrimination) processes of visual perception. This encourages us to extend positive cases (i.e., 128,620 images featuring various colonoscopic findings) with descriptive captions. For this purpose, a straightforward way to create captions is to wrap the category name in a basic template, like “a photo of a [category]” as used by Radford et al. [150]. However, these simple sentences tend to yield suboptimal multimodal alignment, as they are less informative. As shown in Fig. (4-e), we introduce a pipeline to generate more descriptive captions. We interact with GPT-4V [4] using a custom prompt for each colonoscopy image. These prompts act as a prior, conditioned by the image’s category. Compared to simple sentences, the generated captions explain three features. First, our generated captions describe the unique patterns of the target, such as “smooth, round, and nodular appearance”, providing details of the surface of the object. Second, conditioned by category priors, our captions can better differentiate between benign hyperplastic polyps and malignant lesions, describing lesion as “surface appears intact without evidence of erosion or ulceration”. Third, our captions cover not only the lesion itself but also the surroundings, e.g., “mucosa appears slightly erythematous, but without significant inflammatory changes”, offering a more holistic and accurate interpretation.
Figure 5: Response comparison for colonoscopy image classification. We evaluate the zero-shot language responses from three AI chatbots against the response from our multimodal model, ColonGPT.
[…]
5.2 Proposed multimodal language model: ColonGPT
$\bullet$ Motivation. As illustrated in Fig. 5, three AI chatbots – LLaVA-v1.5 [8], LLaVA-Med-v1.5 [280], and GPT-4V [4] – are evaluated for their zero-shot language response capabilities. They often produce inaccurate or vague responses and cannot be readily adaptable to specific colonoscopy tasks. This motivates us to develop a colonoscopy-specific MLM for the community. As a result, ColonGPT classifies the image as “polyp” according to the user instructions, allowing for more precise and personalised applications.
$\bullet$ Overview. We strive to verify the efficacy of language models (LMs) in interpreting both visual and textual signals within the field of medical optical imaging. We present a baseline model, called ColonGPT, to execute colonoscopy tasks following human instructions. As shown in the left of Fig. 6, we follow the framework of MLM [8], which typically involves four basic components. (a) A language tokenizer translates a human instruction $\mathbf{X}_{q}$ into a sequence of tokens $\mathbf{T}_{q}\!\in\!\mathbb{R}^{N_{q}\times D}$, where $N_{q}$ signifies the length of textual tokens and $D$ represents the embedding dimension. (b) A visual encoder, typically based on a transformer, condenses a colonoscopy image $\mathbf{X}_{v}\!\in\!\mathbb{R}^{H\times W\times 3}$, with height $H$ and width $W$, into a flattend visual embedding $\mathbf{E}_{v}\!\in\!\mathbb{R}^{\frac{HW}{P^{2}}\times d}$. Here, $P$ denotes the patch size and $d$ refers to the token dimension. (c) A multimodal adapter transforms the visual embedding $\mathbf{E}_{v}$ into $N_{v}$ numeric tokens $\mathbf{T}_{v}\!\in\!\mathbb{R}^{N_{v}\times D}$, matching the language dimension $D$ with $\mathbf{T}_{q}$. (d) Finally, a language model receives the concatenated visual tokens $\mathbf{T}_{v}$ and text tokens $\mathbf{T}_{q}$ as input. Using the chain rule of probability, a sequence $\mathbf{Y}$ of length $L$ is generated in an autoregressive way, formulated as $p(\mathbf{Y})\!=\!\prod^{L}_{i=1}p(y_{i}\mid\mathbf{T}_{v},\mathbf{T}_{q},\mathbf{Y}_{<i})$, where $\mathbf{Y}_{<i}=[y_{1},y_{2},\cdots,y_{i-1}]$ is the sequence of predicted language tokens indexed before $i$.
Figure 6: Details of our multimodal language model, ColonGPT.
$\bullet$ Multigranularity multimodal adapter. Previous works [7, 8, 9] generally employ a multilayer perceptron architecture as a multimodal adapter, typically consisting of triple linear layers with intervening GELUs. However, handling all visual tokens introduces redundancy because not every token is equally significant, and it also incurs higher computational costs given the quadratic complexity in relation to the number of input tokens. To embrace these challenges, we propose a multimodal adapter that incorporates multigranularity pooling layers between two linear layers. As illustrated in the right of Fig. 6, we transform the embedding $\mathbf{E}_{v}\!\in\!\mathbb{R}^{\frac{HW}{P^{2}}\times d}$ from $d$-dim to $D$-dim using a linear layer followed by a GELU, then reshape it into the spatial format $\mathbf{F}_{v}\!\in\!\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D}$. To reduce the number of visual tokens while avoiding performance drops, we roll out three modifications for the pooling phase, each validated in §5.3.2. (a) Multigranularity views. We add a set of adaptive average pooling operations with $N$ kernel sizes $\{s_{1},\dots,s_{N}\}$ to obtain multigranularity features. In particular, this adaptive operator accommodates input sequences of varying lengths, i.e., the pooled feature for the kernel size $s_{n}$ is shaped as $\mathbb{R}^{s_{n}\times s_{n}\times D}$. (b) Positional encoding. Inspired by [281], we enhance the spatial information within each pooled feature by applying a 2D convolutional layer with the appropriate zero-padding setting. By default, a zero-pixel boundary is added around the input feature. (c) Global view. We also use a global average pooling layer with kernel size $1$ on the feature $\mathbf{F}_{v}$ to obtain a global view with the shape of $\mathbb{R}^{1\times 1\times D}$. Lastly, we reshape each pooled feature into flattened vectors: $\{\mathbf{F}_{p}^{n}\!\in\!\mathbb{R}^{s_{n}^{2}\times D}\}_{n=1}^{N}$ and $\mathbf{F}_{g}\!\in\!\mathbb{R}^{1\times D}$. We concatenate these vectors and process the resulting vector through the second linear layer to produce the final visual tokens $\mathbf{T}_{v}\!\in\!\mathbb{R}^{N_{v}\times D}$, where $N_{v}=(s_{1}^{2}+\cdots+s_{N}^{2}+1)$ denotes the length of the visual tokens.
$\bullet$ Implementation. Our model can be integrated with modern off-the-shelf visual encoders and LMs. To improve reproducibility for average users, we implement ColonGPT in an resource-friendly way. First, we employ SigLIP-SO (0.4B parameters) [5] as the visual encoder, with an input resolution of $H\!=\!W\!=\!384$, a patch size of $P\!=\!14$, and a visual embedding dimension of $d\!=\!1152$. This configuration yields a visual embedding $\mathbf{E}_{v}$ with a shape $\mathbb{R}^{729\times 1152}$, where the number of visual tokens is $729=\lfloor\frac{384}{14}\rfloor^{2}$. In addition, Phi-1.5 (1.3B parameters) [6] serves as the language tokenizer and language model, with a embedding size of $D\!=\!2048$. To reduce computational cost, the size of pooling kernels is set to $\{s_{1},s_{2}\}\!=\!\{14,7\}$, significantly reducing the visual tokens $N_{v}$ from $729$ to $246$, a reduction of $66.26\%$. This design allows us to complete training in five hours, facilitating rapid proof-of-concept development.
$\bullet$ Training recipe. We implement our model using PyTorch library, accelerated by four NVIDIA A100-40GB GPUs. The AdamW optimiser is used with an initial learning rate of 2e-3 and a cosine learning rate scheduler. Our ColonGPT is trained on the combination of training and validation dialogues from ColonINST. The complete training runs for three epochs, with a batch size of 16 per GPU and a gradient accumulation every 2 steps. During training, the visual encoder is frozen, focusing on the trainable multimodal adapter and the LM. For efficiency, we utilise the low-rank adaptation (LoRA [282]) strategy for LM, with a rank of $r\!=\!128$ and a scaling factor of $\alpha\!=\!256$.
TABLE VII: Multimodal benchmark for three conversational tasks. “LoRA” refers to fine-tuning using low-rank adaptation [282]. “EXT” indicates the use of pre-trained weights on extra data. We compare the results on the seen samples from the validation set and the unseen samples from the testing set of ColonINST. The symbol $\uparrow$ signifies that a higher score reflects better performance.
Visual encoder Language model CLS task ($\mathcal{A}\uparrow$) REG task ($\mathcal{A}\uparrow$) REC task (IoU $\uparrow$)
Model (input shape/URL) (model size/URL) No. LoRA EXT seen unseen seen unseen seen unseen
MiniGPT-v2 [7] EVA-G/14 (448px/link) LLaMA2 (7B/link) #A1 #A2 91.49% 90.00% 77.93% 76.82% 94.69% 87.65% 72.05% 70.23% 23.45% 27.97% 15.36% 31.13%
LLaVA-v1 [277] CLIP-L/14 (224px/link) Vicuna-v1.3 (7B/link) #B1 #B2 87.86% 89.61% 72.08% 42.17% 84.55% 86.87% 68.11% 46.85% 20.05% 21.81% 12.72% 3.24%
LLaVA-v1.5 [8] CLIP-L/14 (336px/link) Vicuna-v1.5 (7B/link) #C1 #C2 92.97% 93.33% 79.10% 80.89% 98.58% 99.32% 70.38% 72.88% 55.72% 61.97% 34.32% 42.31%
Bunny-v1.0-3B [9] SigLIP-SO (384px/link) Phi2 (2.7B/link) #D1 #D2 91.16% 92.47% 75.50% 79.50% 96.61% 96.02% 69.45% 75.08% 46.24% 54.00% 31.24% 41.48%
MGM-2B [283] CLIP-L/14 (336px/link) & ConvNeXt-L (768px/link) Gemma (2B/link) #E1 #E2 92.97% 93.24% 78.99% 78.69% 98.17% 98.75% 69.81% 74.30% 39.78% 57.25% 16.00% 25.23%
MobileVLM-1.7B [284] CLIP-L/14 (336px/link) MobileLLaMA (1.4B/link) #F1 #F2 93.02% 93.64% 78.75% 80.44% 97.78% 97.87% 73.14% 78.03% 47.30% 51.36% 31.46% 34.80%
LLaVA-Med-v1.0 [280] CLIP-L/14 (224px/link) LLaMA1 (7B/link) #G1 #G2 93.52% 93.84% 78.04% 77.38% 97.74% 97.35% 75.07% 75.25% 41.60% 39.43% 24.89% 20.85%
LLaVA-Med-v1.5 [280] CLIP-L/14 (224px/link) Mistral-v0.2 (7B/link) #H1 #H2 93.62% 87.22% 79.24% 66.51% 99.30% 90.40% 73.05% 70.00% 64.69% 13.39% 41.97% 12.95%
ColonGPT (Ours) SigLIP-SO (384px/link) Phi1.5 (1.3B/link) - 94.02% 85.81% 99.02% 83.42% 65.89% 45.77%
Figure 7: Illustration of ColonGPT’s multimodal capabilities. Our model can execute various multimodal colonoscopy tasks through conversational interactions, including comprehension (CLS, REG), localisation (REC), and captioning (CAP) based.
5.3 Experiments
5.3.1 Multimodal benchmark
$\bullet$ Model competitors. To establish a widely accepted multimodal benchmark for the community, we select eight popular MLMs as competitors, including six general-purpose and two medically tailored models. As shown in Tab. VII, each competitor has two training setups depending on whether it uses LoRA [282] or initialises knowledge from additional pre-training data. We retrain all competitors using the combined training and validation dialogues from ColonINST, as used in our ColonGPT.
$\bullet$ Evaluation protocols. We quantitatively evaluate three conversational tasks for the multimodal benchmark. For the two classification-based tasks, namely CLS and REG, we adopt the accuracy metric ($\mathcal{A}$) to calculate the ratio of correctly predicted categories to the total number of predictions. For the REC task, we use the intersection over union (IoU) metric to measure the localisation precision. Furthermore, due to the subjective nature of language in the CAP task, we qualitatively analyse the medical accuracy of the responses by verifying the correct identification of the anatomical structures and category names visible in the images, or relevant clinical information descriptions.
$\bullet$ Learning ability. We begin by conducting an open-book test for each model to quantitatively measure how effectively each model has internalised the visual and linguistic patterns from the training phase. Specifically, we evaluate each model on the samples they have seen during training, i.e., 27,606 validation dialogues. The “seen” columns in Tab. VII show that we achieve the highest scores in the CLS ($\mathcal{A}\!=\!94.02\%$) and REC ($\text{IoU}\!=\!65.89\%$) tasks. This implies that ColonGPT has a better learning ability, which allows it to classify images and understand reference expressions related to specific visual regions. Furthermore, we achieve an accuracy of $99.02\%$, below $99.32\%$ achieved by LLaVA-v1.5 (#C2). This gap ($\Delta\!=\!0.3\%$) is due to a 7B-level LM used in LLaVA-v1.5, which has also been pretrained on additional data and thus gained more knowledge.
$\bullet$ Generalisation ability. We further conduct a closed-book test to examine each model’s ability to generalise knowledge to unseen conditions, i.e., 165,830 testing samples of ColonINST. The “unseen” columns in Tab. VII consistently reveal our superior performance in unseen samples in all three tasks. Recall that our model performs worse than LLaVA-v1.5 in the REG task for seen scenarios. However, this gap is overtaken when exposed to unseen scenarios, where we achieve an accuracy of $83.42\%$, even exceeding the 7B-level LLaVA-v1.5 (#C2) by $10.54\%$. Compared to the runner-up model, LLaVA-Med-v1.5 (#H1), our model surpasses it by $1.2\%$ in the seen scenarios for the REC task, and this gap is further widened by $3.8\%$ in the unseen scenarios. These results show the potential of ColonGPT as a conversational assistant for multimodal colonoscopy tasks, especially in generalising to unseen data.
TABLE VIII: Diagnostic studies of three core components in ColonGPT.*”: interpolate the position embeddings for higher resolution, specifically from 224px to 384px. Our default configurations are shaded with a gray background.
(a) Different presentations from visual encoder
Visual encoder input/URL CLS REG REC
ConvNeXtV2-L 384px/link 82.95% 78.63% 33.74%
ViT-L 384px/link 82.16% 77.04% 40.78%
MAE-L* 384px/link 80.85% 75.87% 38.53%
MAE-L 224px/link 81.95% 77.62% 43.25%
DINOv2-L* 384px/link 35.03% 22.91% 6.79%
DINOv2-L 224px/link 21.22% 7.96% 2.69%
CLIP-L 336px/link 83.99% 78.67% 41.54%
SigLIP-SO 384px/link 85.81% 83.42% 45.77%
(b) Multigranuarity multimodal adapter
token (ratio) CLS REG REC
MLP baseline 729 (100.00%) 83.53% 81.80% 43.70%
$\{16,8,1\}$ 321 (44.03%) 84.39% 80.90% 46.37%
$\{14,7,1\}$ 246 (33.74%) 85.81% 83.42% 45.77%
$\{14,7\}$ 245 (33.61%) 85.01% 82.49% 43.62%
$\{12,6,1\}$ 181 (24.83%) 83.74% 81.60% 45.94%
$\{10,5,1\}$ 126 (17.28%) 84.28% 82.01% 46.46%
$\{8,4,1\}$ 81 (11.11%) 84.70% 81.36% 45.30%
w/o Pos. Enc. 246 (33.74%) 84.50% 82.91% 40.09%
(c) Fine-tuning strategy
Strategy $r$ $\alpha$ CLS REG REC
full-tuning - - 78.06% 73.79% 50.20%
LoRA 4 8 85.43% 82.75% 45.02%
LoRA 8 16 84.45% 80.78% 44.98%
LoRA 16 32 84.39% 80.81% 45.90%
LoRA 32 64 84.91% 82.73% 45.56%
LoRA 64 128 83.84% 81.19% 43.57%
LoRA 128 256 85.81% 83.42% 45.77%
LoRA 256 512 82.93% 79.96% 48.27%
$\bullet$ Qualitative analysis. Fig. 7 illustrates our model’s three multimodal abilities across four conversational tasks. (a) Comprehension ability: In the CLS task, we identify subtle visual features, distinguishing “high grade dysplasia” from “adenoma” in visually similar images. In the REG task, we correctly translate complex visual features from the given coordinates into precise medical terminology. (b) Localisation ability: This entails ColonGPT understanding language query and localising visual target within a complex colon environment. The outputs of the REC task showcase ColonGPT’s precision in localising specified expressions using bounding boxes. (c) Captioning ability: This requires the synthesis of visual information into coherent, clinically relevant text. Our model provides descriptions of a pedunculated polyp, detailing its morphology, contextual characteristics, and potential clinical relevance. Additionally, ColonGPT can describe the treatment procedure when an instrument is present, e.g., “The polyp is being examined using a snare.”
5.3.2 Diagnostic studies
$\bullet$ Visual encoder. Our diagnostic experiment begins with an inquiry – What types of visual representations are appropriate for multimodal colonoscopy data? We prepare four sets of representations from various large-scale pre-training strategies: supervised learning (ConvNeXtV2 [285], ViT [100]), reconstructive learning (MAE [286]), and contrastive learning using vision-only data (DINOv2 [287]) or VL data (CLIP [288], SigLIP [5]). As shown in Tab. (8-a), all encoders use pre-trained weights from Huggingface. To ensure consistency, we manually interpolate the smaller position embedding for MAE and DINOv2 from 224px to 384px (marked with *), while leaving the default input of the remaining models unchanged. Our observation reveals that contrastive learning encoders using VL data outperform other strategies. This suggests that visual representations pre-aligned with weak texts during their pre-training facilitate us to transform visual embeddings into the language space. Regarding the other unimodal encoders, both supervised learning methods (ConvNeXtV2, ViT) and the reconstructive approach (MAE) give satisfactory feedback. However, the vision-only contrastive learning model (DINOv2) struggles, especially in the REC and REG tasks. Multiple efforts to optimise DINOv2 with various hyperparameters, including input size, model size, and learning rate, did not produce acceptable results for multimodal colonoscopy tasks.
$\bullet$ Multigranularity multimodal adapter. It serves as a key component in linking vision and language modalities, reducing visual tokens to mitigate computational overhead. As detailed in Tab. (8-b), we analyse its effectiveness from three perspectives. (a) How to configure the pooling kernels? As a reference, we initialise a baseline variant of ColonGPT with a multimodal adapter used in [8, 7], which employs a pure MLP architecture to process all input tokens equally. We then gradually decrease the size of the pooling kernels across five variants: $\{16,8,1\}$, $\{14,7,1\}$, $\{12,6,1\}$, $\{10,5,1\}$, and $\{8,4,1\}$. Concerning the performance-cost trade-offs, our setup $\{14,7,1\}$ is optimal. It decreases the visual tokens from 100% (729 tokens) to 33.74% (246 tokens) while maintaining impressive results across three conversational tasks. To illustrate, we observe performance gains of 2.28%, 1.62%, and 2.07% in the CLS, REG, and REC tasks, respectively. (b) Is global context necessary? We remove the global view from our default setup $\{14,7,1\}$ for the multigranularity adapter, producing a controlled variant with setup $\{14,7\}$. The performance then declines, indicating the necessity of capturing the global context within visual embeddings for improved outcomes. (c) Is positional encoding important? As shown in the last row of Tab. (8-b), our setup without positional encoding shows a significant performance drop in REC task, from 45.77% to 40.09%. This suggests that the relative position information for the visual sequence is crucial for the localisation task.
$\bullet$ Fine-tuning strategy. Lastly, we investigate how to effectively tune our model on multimodal colonoscopy data. As shown in Tab. (8-c), we initiate a set of variants to tune the language model, Phi1.5. It includes seven variants with different setups for LoRA and a full-tuning variant as a reference point. The best performance was observed in the LoRA variant with configuration $r/\alpha=128/256$, achieving 85.81% and 83.42% in the CLS and REG tasks, respectively. In addition, the other two variants achieve outstanding scores in the REC task, with the full-tuning variant achieving 50.20% and the LoRA with configuration $r/\alpha\!=\!256/512$ reaching 48.27%. This implies that more tunable parameters for the language model could allow ColonGPT to capture more complex patterns in the spatial planes. Therefore, we believe that there remains room to improve the localisation ability of ColonGPT in the REC task, such as increasing the rank factor $r$ and finding a matched scaling factor $\alpha$. For training efficiency, we finally choose $r/\alpha\!=\!128/256$ as our default configuration, as it allows us to reduce the training time from eight to roughly five hours.
5.3.3 Empirical takeaways
This study represents a preliminary exploration of multimodal instruction tuning techniques in colonoscopy. We unify the multimodal and multitask paradigms in a causal language model, which features two insights: interpreting visual content within the linguistic space and tackling various visual tasks under a next-token prediction framework. We finally derive lessons from experiments that may guide future advances in multimodal research.
$\bullet$ Embracing data scarcity. In general, MLMs [277, 8] opt for a two-stage strategy trained on massive data, e.g., $\sim$558K samples for multimodal alignment, followed by $\sim$665K instruction tuning samples to ensure human compliance. Alternatively, we adopt a single-stage strategy to directly fine-tune ColonGPT on comparatively smaller training data with $\sim$285K instructions. This strategy appears to be effective in colonoscopy, a data-limited scenario. We suggest two feasible ways to compensate for this data-centric issue. (a) Scaling up data size is a straightforward way to improve the domain-specific representation ability. A cost-efficient way is to consider synthesised data once the public data sources are used up [289]. (b) Diversifying the human-machine dialogue can efficiently train an AI specialist for colonoscopy applications. This involves expanding question-answer pairs with advanced AI chatbots and organising more executable tasks, such as converting masks into polygons for segmentation [290] or modelling multiframe correlations for video analysis [291].
$\bullet$ Efficiency drives progress. As discussed above, we take less data to obtain greater performance than other model rivals. This success benefits from the way we build ColonGPT. (a) Colonoscopy data inherently contains redundant information, such as the fact that most mucosal surfaces are similar, as well as camouflaged patterns between benign lesions and their surroundings, as discussed in §2.2. To reduce redundancy, we propose a multigranularity multimodal adapter that selectively samples tokens without compromising performance. For future improvement, we can draw on the wisdom of previous token reduction techniques [292]. (b) As shown in Tab. VII, the Phi1.5 model [6], although lightweight, shows surprising efficiency, even outperforming other 7B-level competitors. This indicates that larger models appear to require more colonoscopy data. Thus, future efforts should prioritise enhancing training and inference efficiency, especially for the medical field, rather than racing with massive computational resources. A promising idea to streamline the MLM framework using an encoder-free solution [293] to interpret visual pixels.
$\bullet$ Improving spatial perception. We observe that the ability to accurately locate targets given descriptions remains limited. This is evident in the REC results shown in Tab. VII, where IoU scores fall below 50% in most models when tested on unseen samples. To break through this performance bottleneck, we suggest two potential routes. (a) In constructing ColonGPT, we leverage a pre-trained visual encoder and a language model from the general domain. This approach presents challenges of the gaps between general and medical optical data, as well as the gap between vision and language modalities. As recommended in [294], pre-training and pre-aligning the multimodal space before instruction tuning would be a promising approach to alleviate these issues. (b) The next-token prediction framework of causal language models may struggle with arithmetic tasks due to the snowballing error resulting from the chain rule [295]. For example, LMs are not responsible for accurately predicting coordinates in the REC task. We encourage that the vision and language parts of the next-generation framework can handle their respective roles, such as a parallel framework [296] that predicts segmentation masks and generates language captions, simultaneously.
6 Conclusion
We investigate the frontiers in intelligent colonoscopy techniques and their broader implications in the multimodal field. Our structure follows two primary threads. First, we survey the landscape of four colonoscopic scene perception tasks and sort out the key challenges and understudied areas. Second, our survey reveals that multimodal research in colonoscopy is underexplored. To embrace this, we contribute three initiatives to the community: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-specific multimodal language model ColonGPT, and a multimodal benchmark. Importantly, multimodal colonoscopy research is rapidly advancing. Future developments can build on recent breakthroughs, such as executing role-playing tasks with agents [297].
Acknowledgments
This research was supported by NSFC (NO.62476143). We express our sincere gratitude to Yu-Cheng Chou (JHU) and Stephen Gould (ANU) for interesting discussions.

### v2

@@
 With this goal, we begin by assessing the current data-centric and model-centric landscapes through four tasks for colonoscopic scene perception, including classification, detection, segmentation, and vision-language understanding.
-This assessment enables us to identify domain-specific challenges and reveals that multimodal research in colonoscopy remains open for further exploration.
-To embrace the coming multimodal era, we establish three foundational initiatives: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-designed multimodal language model ColonGPT, and a multimodal benchmark.
-To facilitate ongoing monitoring of this rapidly evolving field, we provide a public website for the latest updates: https://github.com/ai4colonoscopy/IntelliScope.
+Our assessment reveals domain-specific challenges and underscores the need for further multimodal research in colonoscopy.
+To address these gaps, we establish three foundational initiatives: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-designed multimodal language model ColonGPT, and a multimodal benchmark.
+To facilitate continuous advancements in this rapidly evolving field, we provide a public website for the latest updates: https://github.com/ai4colonoscopy/IntelliScope.
 Index Terms:
-Colonoscopy, Survey, Vision-language, Multimodal Language Model, Medical Image, Abdomen, Healthcare AI.
-1 Introduction
+Colonoscopy, Survey, Vision-language, Multimodal language model, Multimodal benchmark, Abdomen, Healthcare AI.
+I Introduction
 Despite declining colorectal cancer (CRC) rates in high-income countries, it remains the third most diagnosed cancer worldwide and is increasing in developing countries [1].
-Colonoscopy, as an efficient method for CRC screening, utilises a flexible camera-equipped tube to visually examine the colon’s interior.
+As an efficient method for CRC screening, colonoscopy utilises a flexible camera-equipped tube to visually examine the colon’s interior.
 As illustrated in Fig.
@@
 This success motivates us to investigate the frontiers in intelligent colonoscopy.
-Colonoscopy, an endoscopic optical imaging technique, usually presents visual patterns (e.g., non-uniform illumination, homogeneity) that differ from those of general-purpose imaging data, e.g., ImageNet [3], due to the complex and folded anatomy of the colon.
+Colonoscopy, an endoscopic optical imaging technique, usually presents challenges such as non-uniform illumination and homogeneity of visual patterns, that differ from those of general-purpose imaging data, e.g., ImageNet [3], due to the complex and folded anatomy of the colon.
 This suggests that special methods are needed to interpret the colonoscopic data.
@@
 Our analysis reveals that multimodal research in colonoscopy remains largely untapped.
-To bridge this gap, we contribute three efforts to the community, as illustrated in Fig.
+To bridge this gap, we contribute the following three efforts to the community, as illustrated in Fig.
 (1-b).
-Figure 1: Introductory diagram.
-We depict (a) the anatomy of the large intestine (colon) within the digestive tract, the polypectomy procedure during colonoscopy examination, and the components of a colonoscope.
-The bottom figure (b) summarises three highlights of this study.
+Figure 1: Overview of colonoscopy and study highlights: We depict (a) the anatomy of the large intestine (colon) within the digestive tract, the polypectomy procedure during colonoscopy examination, and the colonoscope components.
+(b) summarises three highlights of this study.
 Figure 2: Colonscopic scene perception from visual to multimodal perspectives.
@@
 2) from both data-centric and model-centric perspectives.
-Our investigation summarises key features of 63 datasets and 137 representative deep techniques published since 2015.
+Our investigation summarises key features of 63 datasets and 137 representative deep learning techniques published since 2015.
 Additionally, we highlight emerging trends and opportunities for future study.
@@
 To ensure reproducibility for average community users, we implement ColonGPT in a resource-friendly way, using a 0.4B-parameter visual encoder SigLIP-SO [5] and a 1.3B-parameter lightweight language model Phi1.5 [6].
-Unlike previous vision-language (VL) linking methods [7, 8, 9] that employ multilayer perceptrons to handle equally all tokens from the visual encoder, we propose a multigranularity adapter to selectively sample visual tokens according to their significance.
-This strategy reduces the visual tokens to $\sim$34of the original number without compromising performance, securing the top spot in our newly-created multimodal benchmark across three tasks.
-Importantly, our ColonGPT can be trained within five hours on four A100-40GB GPUs, facilitating rapid proof-of-concept development for subsequent research.
+Unlike previous vision-language (VL) linking methods [7, 8, 9] that employ multilayer perceptrons to handle all tokens equally from the visual encoder, we propose a multigranularity adapter to selectively sample visual tokens according to their significance.
+This strategy reduces the visual tokens to $\sim$34of the original number without compromising performance, securing the top spot in our newly created multimodal benchmark across three tasks.
+Importantly, our ColonGPT can be trained within seven hours on two NVIDIA H200 GPUs, facilitating rapid proof-of-concept development for subsequent research.
 Scope.
@@
 Earlier surveys on traditional [10] and deep learning [11, 12, 13, 14] methods conducted before 2020 are now out of date.
-Although a recent study [15] explores various applications of colonoscopy, such as quality analysis and abnormality detection, it lacks numerical validation.
+Although a recent study [15] explores various applications of colonoscopy, such as quality analysis and anomaly detection, it lacks numerical validation.
 Other benchmarks [16, 17, 18] are limited to specific narrow research subfields.
-By contrast, we delve into four tasks for colonoscopic scene perception and evaluate their current state to sort out key challenges and under-researched areas.
-Importantly, our vision goes beyond by laying the foundations for the coming era, a multimodal world.
-To embrace this era, we further undertake three initiatives: a multimodal instruction tuning dataset, a multimodal language model, and a multimodal benchmark for the community.
+By contrast, we delve into four tasks related to colonoscopic scene perception and evaluate their current state to identify key challenges and under-researched areas.
+Importantly, our vision goes beyond by laying the foundations for the upcoming advancements in the multimodal domain.
+To this end, we undertake three initiatives to community: a multimodal instruction tuning dataset, a multimodal language model, and a multimodal benchmark.
 Organisation.
-The remaining sections are structured as such: §2 provides a historical background and discusses the domain-unique challenges.
-§3 investigates 63 colonoscopy related datasets, followed by a survey of 137 deep models in §4.
-In §5, we introduce three initiatives towards the multimodal era: the creation of ColonINST, the technical details of ColonGPT, and a comparative multimodal benchmark along with ablative analyses.
-Finally, this paper is concluded in §6.
+The remaining sections are structured as follows: §II provides a historical background and discusses the unique challenges of the domain.
+In §III, we investigate 63 datasets related to colonoscopy, followed by a survey of 137 deep models in §IV.
+In §V, we introduce three initiatives towards the multimodal era: the creation of ColonINST dataset, the technical details of ColonGPT model, and a comparative multimodal benchmark along with ablative analyses.
+Finally, this paper is concluded in §VI.
 TABLE I: Data statistics for colonoscopy datasets.
@@
   • **NOTE – The MedFMC dataset [70] comprises 23,349 medical images across five modalities. This table only enumerates the categories specific to the endoscopic modality.
-2 Background
-2.1 Origin and evolution
+II Background
+II-A Origin and evolution
 The history of colonoscopy has two key milestones.
@@
 The 21st century brings the AI era, where computer-aided diagnosis systems provide greater precision and reliability in procedures [80].
-This study explores the transformative impact of intelligent techniques for colonoscopy, which is a type of endoscopy [81], while other related techniques such as laparoscopy [82] are summarised in our appendix.
-2.2 Intrinsic traits and domain-unique challenges
+This study explores the transformative impact of intelligent techniques for colonoscopy, a type of endoscopy [81], while other related areas such as laparoscopy [82] are discussed in appendix.
+II-B Intrinsic traits and domain-unique challenges
 We summarise five unique challenges associated with colonoscopic vision tasks, primarily caused by procedural aspects and imaging conditions during a colonoscopy.
@@
 These issues require a robust response from AI models to inherent morphological and colour fluctuations.
-3 Revisiting Colonoscopy Data
-3.1 Medical data for colonoscopy
+III Revisiting Colonoscopy Data
+III-A Medical data for colonoscopy
 Tab.
@@
 $\bullet$ Detection datasets provide both categorical and localisation labels for targets of interest, such as colonic diseases [54, 57, 62], accessory instruments [42, 34, 42], polyps [48, 59, 23, 64, 72], endoscopic artefacts [45, 36], and gastrointestinal diseases [56, 43].
-In addition, the organisation of competitions has accelerated growth within the colonoscopy community by establishing shared platforms for data collection and model evaluation, significantly advancing research areas such as the detection of intestinal disease [40] and polyp [61, 69, 38].
+In addition, the organisation of competitions has accelerated growth within the colonoscopy community by establishing shared platforms for data collection and model evaluation, greatly advancing research areas like the detection of intestinal disease [40] and polyp [61, 69, 38].
 $\bullet$ Segmentation datasets for colonoscopy research originate from two sources.
@@
 Kvasir-VQA [74] collects 6,500 question-answer pairs from existing datasets [43, 50], for gastrointestinal diagnostics, such as image captioning, visual question answering.
-3.2 Discussion
+III-B Discussion
 Based on the 63 revisited datasets, we have several data-centric observations that could inspire more future ideas.
@@
 Several data-centric areas remain under-explored, such as temporal lesion evolution recording, granularity improvement, graded severity tagging, and instance-level target annotation.
-(b) Furthermore, label orthogonality, an often overlooked issue, treats labels as isolated entities.
+(b) Label orthogonality, an often overlooked issue, treats classes as isolated entities.
 Current works seldom discuss potential inter-class correlations, such as the co-occurrence of inflammatory bowel disease with erosion symptoms, Crohn’s disease with fistula complications, or colorectal cancer accompanied by bleeding.
@@
 Dataset: CU=CU-ColonDB [25], CDS=ColonoscopicDS [26], Private= private data, HK=HyperKvasir [43], KC=Kvasir-Capsule [56].
-Backbone: CaffeNet [97], D-121=DenseNet121 [98], R-12/-18/-50/-101=ResNet12/18/50/101 [99], ViT-S16 or ViT-B16 [100], MobV2=MobileNetV2 [101], R50-Att=ResNet50 with attention module [102], C3D [103], Inc-v3=Inceptionv3 [104], I3D [105].
+Backbone: CaffeNet [97], D-121=DenseNet121 [98], R-12/-18/-50/ -101=ResNet12/18/50/101 [99], ViT-S16/ViT-B16 [100], MobV2=MobileNetV2 [101], R50-Att=ResNet50 with attention module [102], C3D [103], Inc-v3=Inceptionv3 [104], I3D [105].
 “Customised” means a base network modified for the current task or a model independent of the base network choice.
@@
 These side outputs typically receive supervision from additional supervisory signals, such as boundary cues.
-4 Revisiting Colonoscopy Models
+IV Revisiting Colonoscopy Models
 In this section, we investigate 137 deep learning models for colonoscopic scene perception, sourced from leading conferences or journals published since 2015.
-First, three widely recognised topics are described, including 18 classification models (§4.1), 24 detection models (§4.2), and 86 segmentation models (§4.3).
+First, three widely recognised topics are described, including 18 classification models (§IV-A), 24 detection models (§IV-B), and 86 segmentation models (§IV-C).
 Their architectural designs are classified into five subtypes as presented in Fig.
 3.
-Lastly, we discuss nine VL-related models in §4.4.
-4.1 Classification models
+Lastly, we discuss nine VL-related models in §IV-D.
+IV-A Classification models
 $\bullet$ Input phase.
@@
 Intrator et al. [147] MICCAI’23 self-supervised polyp re-identification Private Private R-50v2 OS MF#2 US -
-V2I-DETR [148] arXiv’24 video-to-image knowledge distillation SUN SUN R-50 OS BF#2 FS -
+V2I-DETR [148] BIBM’24 video-to-image knowledge distillation SUN SUN R-50 OS BF#2 FS -
 $\bullet$ Output phase.
@@
 The Kvasir series [28, 43, 56] offers valuable sources for further study.
-We will explore these public data on their potential synergy in §5.1.
+We will explore these public data on their potential synergy in §V-A.
 (c) Several new task settings have emerged in colonoscopy.
@@
 DAFON [115] solve an open-set classification problem within a few-shot framework.
-4.2 Detection models
+IV-B Detection models
 $\bullet$ Input phase.
@@
 Gao et al.[235] TMI’24 in-context learning, dual similarity checkup C3 C3 SAM BF#2 - WS
-U-KAN [236] arXiv’24 U-shaped Kolmogorov-Arnold network [237] C6 C6 Customised MF#1 - FS Link
 SliceMamba [236] arXiv’24 vision state space model, bidirectional slice scan C6, KS C6, KS Customised MF#1 - FS -
@@
 SAM2-UNet [238] arXiv’24 SAM2 [166], adapter-based tuning C6, KS C6, ES, KS, C3, ETIS SAM2 BF#1 EX FS Link
+U-KAN [239] AAAI’25 U-shaped Kolmogorov-Arnold network [240] C6 C6 Customised MF#1 - FS Link
 Video-based models Puyal et al. [241] MICCAI’20 temporal correlation via hybrid 2D/3D CNNs Private, KS Private R-101 MF#2 - FS -
@@
 TABLE V: Comparison of image polyp segmentation models.
-They are evaluated using the mean scores (%) of structure measure ($\mathcal{S}$ [258]) and Dice coefficient ($\mathcal{D}$) on two test sets, with boxplots illustrating the distribution of their consistency and variability across test cases.
+Models are evaluated using the mean scores (%) of structure measure ($\mathcal{S}$ [258]) and Dice coefficient ($\mathcal{D}$) on two test sets, with boxplots illustrating the distribution of their consistency and variability across test cases.
+The rankings for each model are marked in grey colour.
 Kvasir-SEG (100 test images) [38] CVC-ClinicDB (62 test images) [21]
@@
 SFA [172] 78.14 [#14] 72.31 [#14] 79.33 [#14] 70.06 [#14]
-4.3 Segmentation models
-Compared to the above two topics, the segmentation research appears to be well established. This is evidenced by the extensive amount of research documented in Tab. IV.
+IV-C Segmentation models
+Compared to the above two topics, the segmentation research appears to be well-established (see Tab. IV).
 $\bullet$ Input phase.
@@
 Moreover, some methods adopt a hybrid strategy, e.g., both the subtraction operator and the edge-aware loss are used in MSNet [184].
+[…]
+Figure 4: Details of ColonINST.
+(a) Three sequential steps to create the instruction tuning dataset for multimodal research.
+(b) Numbers of colonoscopy images designated for training, validation, and testing purposes.
+(c) Data taxonomy of three-level categories.
+(d) A word cloud of the category distribution by name size.
+(e) Caption generation pipeline using the VL prompting mode of GPT-4V [4].
+(f) Numbers of human-machine dialogues created for four tasks.
 $\bullet$ Output phase.
@@
 $\bullet$ Remarks.
-To reflect current field progress, we analyse 14 open-source image segmentation models on two popular test datasets, as shown in Tab.
+To reflect current field progress, we analysed 14 open-source image segmentation models on two popular test datasets, as shown in Tab.
 V.
 All models are trained on the same dataset from Fan et al.’s benchmark [175].
-First, we observe that current learning strategies are underexplored, as evidenced by BoxPolyp, a weakly supervised model obtaining the best $\mathcal{D}$ score (91.84%) on Kvasir-SEG.
+First, we observe that current learning strategies are underexplored, as evidenced by BoxPolyp, a weakly supervised model that obtains the best $\mathcal{D}$ score (91.84%) on Kvasir-SEG.
 In addition, some models achieve better performance, yet exhibit wider interquartile ranges in boxplots, indicating their variability in predictions.
@@
 From these results, we suggest several promising opportunities for future study.
-(a) The current gold benchmark [175] comprises less than 1.5K samples and is focused on one category (polyp).
+(a) The current gold benchmark [175] comprises less than 1.5K samples and is focused on a category (polyp).
 In general, scaling up both the data size and diversity could be a natural way to improve robustness and generalisability.
@@
 We can adopt innovations from similar topics [263].
-[…]
-Figure 4: Details of the established ColonINST.
-(a) Three sequential steps to create the instruction tuning dataset for multimodal research.
-(b) Numbers of colonoscopy images designated for training, validation, and testing purposes.
-(c) Data taxonomy of three-level categories.
-(d) A word cloud of the category distribution by name size.
-(e) Caption generation pipeline using the VL prompting mode of GPT-4V [4].
-(f) Numbers of human-machine dialogues created for four downstream tasks.
-4.4 VL models
+IV-D VL models
 Compared to the above three topics, multimodal research has relatively fewer references.
@@
 Future insights can be learnt from existing ideas.
-First, crawling unlabelled image-text data from social media [275] and scientific literature [276] can be used to build domain-specific foundation models.
+First, crawling unlabelled image-text data from social media [275] or the scientific literature [276] can be used to build domain-specific foundation models.
 Second, language models such as GPT-4V [4] can generate diverse professional descriptions, offering an economical and scalable way to expand the knowledge space of the data.
@@
 These techniques opt for a decoder-only strategy that unifies multiple tasks (e.g., detection, captioning) into a unified autoregressive framework (i.e., next-token prediction).
-These models are flexibly capable of processing input and output texts of varying lengths, without requiring additional task-specific prediction heads for different tasks.
-5 Stepping into the Multimodal Land
+These models are flexibly capable of processing input and output texts of varying lengths, without requiring additional task-specific heads for different tasks.
+V Advancing Multimodal Learning for Colonoscopy
 TABLE VI: Details of instruction tuning dataset ColonINST.
@@
 Instruction tuning [278] is key to instructing the MLMs to execute domain-specific tasks aligned with user preferences.
-This section introduces three initiatives to advance multimodal research: how we create a large-scale instruction tuning dataset, ColonINST (§5.1) and how we train a colonoscopy-specific MLM, ColonGPT (§5.2).
-Finally, as in §5.3, we contribute a multimodal benchmark for conversational colonoscopy tasks, conduct diagnostic studies on ColonGPT, and share lessons from empirical observations.
-5.1 Established instruction tuning dataset: ColonINST
-$\bullet$ Overview.
+This section introduces three initiatives to advance multimodal research: how we create a large-scale instruction tuning dataset, ColonINST (§V-A) and how we train a colonoscopy-specific MLM, ColonGPT (§V-B).
+Finally, we contribute the first multimodal benchmark for conversational tasks (§V-C), conduct diagnostic studies on ColonGPT (§V-D), and share lessons from our empirical observations (§V-E).
+V-A Established instruction tuning dataset: ColonINST
 Fig.
@@
 We intentionally keep these negative samples, as they may be valuable for future gastrointestinal research.
-Fig.
-(4-d) presents a word cloud distribution for all categories names within ColonINST.
+As shown in Fig.
+(4-d), we present a word cloud distribution for all categories names within ColonINST.
 $\bullet$ Caption generation.
@@
 Third, our captions cover not only the lesion itself but also the surroundings, e.g., “mucosa appears slightly erythematous, but without significant inflammatory changes”, offering a more holistic and accurate interpretation.
+[…]
+V-B Proposed multimodal language model: ColonGPT
 Figure 5: Response comparison for colonoscopy image classification. We evaluate the zero-shot language responses from three AI chatbots against the response from our multimodal model, ColonGPT.
-[…]
-5.2 Proposed multimodal language model: ColonGPT
 $\bullet$ Motivation.
@@
 Using the chain rule of probability, a sequence $\mathbf{Y}$ of length $L$ is generated in an autoregressive way, formulated as $p(\mathbf{Y})\!=\!\prod^{L}_{i=1}p(y_{i}\mid\mathbf{T}_{v},\mathbf{T}_{q},\mathbf{Y}_{<i})$, where $\mathbf{Y}_{<i}=[y_{1},y_{2},\cdots,y_{i-1}]$ is the sequence of predicted language tokens indexed before $i$.
-Figure 6: Details of our multimodal language model, ColonGPT.
 $\bullet$ Multigranularity multimodal adapter.
@@
 6, we transform the embedding $\mathbf{E}_{v}\!\in\!\mathbb{R}^{\frac{HW}{P^{2}}\times d}$ from $d$-dim to $D$-dim using a linear layer followed by a GELU, then reshape it into the spatial format $\mathbf{F}_{v}\!\in\!\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D}$.
-To reduce the number of visual tokens while avoiding performance drops, we roll out three modifications for the pooling phase, each validated in §5.3.2.
+To reduce the number of visual tokens while avoiding performance drops, we roll out three modifications for the pooling phase, each validated in §V-D.
 (a) Multigranularity views.
@@
 Lastly, we reshape each pooled feature into flattened vectors: $\{\mathbf{F}_{p}^{n}\!\in\!\mathbb{R}^{s_{n}^{2}\times D}\}_{n=1}^{N}$ and $\mathbf{F}_{g}\!\in\!\mathbb{R}^{1\times D}$.
-We concatenate these vectors and process the resulting vector through the second linear layer to produce the final visual tokens $\mathbf{T}_{v}\!\in\!\mathbb{R}^{N_{v}\times D}$, where $N_{v}=(s_{1}^{2}+\cdots+s_{N}^{2}+1)$ denotes the length of the visual tokens.
+We concatenate these vectors and process the resulting vector through the second linear layer to produce the final visual tokens $\mathbf{T}_{v}\!\in\!\mathbb{R}^{N_{v}\times D}$, where $N_{v}=(s_{1}^{2}+\cdots+s_{N}^{2}+1)$ is the length of visual tokens.
 $\bullet$ Implementation.
 Our model can be integrated with modern off-the-shelf visual encoders and LMs.
-To improve reproducibility for average users, we implement ColonGPT in an resource-friendly way.
+To improve reproducibility for average users, we implement ColonGPT in a resource-friendly way.
 First, we employ SigLIP-SO (0.4B parameters) [5] as the visual encoder, with an input resolution of $H\!=\!W\!=\!384$, a patch size of $P\!=\!14$, and a visual embedding dimension of $d\!=\!1152$.
@@
 To reduce computational cost, the size of pooling kernels is set to $\{s_{1},s_{2}\}\!=\!\{14,7\}$, significantly reducing the visual tokens $N_{v}$ from $729$ to $246$, a reduction of $66.26\%$.
-This design allows us to complete training in five hours, facilitating rapid proof-of-concept development.
-$\bullet$ Training recipe.
-We implement our model using PyTorch library, accelerated by four NVIDIA A100-40GB GPUs.
-The AdamW optimiser is used with an initial learning rate of 2e-3 and a cosine learning rate scheduler.
-Our ColonGPT is trained on the combination of training and validation dialogues from ColonINST.
-The complete training runs for three epochs, with a batch size of 16 per GPU and a gradient accumulation every 2 steps.
-During training, the visual encoder is frozen, focusing on the trainable multimodal adapter and the LM.
-For efficiency, we utilise the low-rank adaptation (LoRA [282]) strategy for LM, with a rank of $r\!=\!128$ and a scaling factor of $\alpha\!=\!256$.
+This design allows us to complete the training in seven hours, facilitating rapid proof-of-concept development.
+Figure 6: Details of our multimodal language model, ColonGPT.
 TABLE VII: Multimodal benchmark for three conversational tasks.
 “LoRA” refers to fine-tuning using low-rank adaptation [282].
-“EXT” indicates the use of pre-trained weights on extra data.
+“EXT” indicates the use of pre-trained weights on extra general data.
 We compare the results on the seen samples from the validation set and the unseen samples from the testing set of ColonINST.
-The symbol $\uparrow$ signifies that a higher score reflects better performance.
+Symbol $\uparrow$ means that a higher score reflects better performance.
 Visual encoder Language model CLS task ($\mathcal{A}\uparrow$) REG task ($\mathcal{A}\uparrow$) REC task (IoU $\uparrow$)
@@
 LLaVA-Med-v1.5 [280] CLIP-L/14 (224px/link) Mistral-v0.2 (7B/link) #H1 #H2 93.62% 87.22% 79.24% 66.51% 99.30% 90.40% 73.05% 70.00% 64.69% 13.39% 41.97% 12.95%
-ColonGPT (Ours) SigLIP-SO (384px/link) Phi1.5 (1.3B/link) - 94.02% 85.81% 99.02% 83.42% 65.89% 45.77%
-Figure 7: Illustration of ColonGPT’s multimodal capabilities.
-Our model can execute various multimodal colonoscopy tasks through conversational interactions, including comprehension (CLS, REG), localisation (REC), and captioning (CAP) based.
-5.3 Experiments
-5.3.1 Multimodal benchmark
+ColonGPT (Ours) SigLIP-SO (384px/link) Phi1.5 (1.3B/link) - 94.06% 83.24% 99.96% 80.18% 85.74% 56.24%
+$\bullet$ Training recipe.
+We implement our model using the PyTorch library, accelerated by two NVIDIA H200 GPUs.
+The AdamW optimiser is used with an initial learning rate of 2e-3 and a cosine learning rate scheduler.
+Our reproducible process consists of two training stages.
+(a) Pre-alignment stage: First, we train the multigranularity multimodal adapter on $\sim$83K image-caption pairs to pre-align the visual encoder with the language model, enabling it to interpret visual tokens.
+At this stage, both the visual encoder and the LM remain frozen, while the focus is on the multimodal adapter with a learning rate of 2e-4.
+(b) Supervised fine-tuning stage: Next, we adapt the model to address colonoscopy-related tasks by fine-tuning with CLS, REG, and REC data, amounting to $\sim$202K image-text pairs.
+For efficiency, the low-rank adaptation strategy (LoRA [282]) is applied to the LM, with a rank of $r\!=\!128$ and a scaling factor of $\alpha\!=\!256$.
+During this stage, we train the adapter and LM with a learning rate of 2e-3 and 2e-4, respectively.
+Each stage lasts three epochs with a batch size of 16 per GPU and gradient accumulation every two steps.
+The duration of training takes about 1.4 hours for pre-alignment and 5.3 hours for fine-tuning.
+V-C Multimodal benchmark
 $\bullet$ Model competitors.
@@
 As shown in Tab.
-VII, each competitor has two training setups depending on whether it uses LoRA [282] or initialises knowledge from additional pre-training data.
-We retrain all competitors using the combined training and validation dialogues from ColonINST, as used in our ColonGPT.
+VII, each competitor has two training setups depending on whether it uses LoRA [282] or initialises knowledge from weights pre-trained on additional general data.
+We retrain each competitor using all the training and validation dialogues from ColonINST.
 $\bullet$ Evaluation protocols.
@@
 For the REC task, we use the intersection over union (IoU) metric to measure the localisation precision.
-Furthermore, due to the subjective nature of language in the CAP task, we qualitatively analyse the medical accuracy of the responses by verifying the correct identification of the anatomical structures and category names visible in the images, or relevant clinical information descriptions.
+Furthermore, due to the subjective nature of language in the CAP task, we qualitatively analyse the medical accuracy of the responses by verifying the correct identification of the anatomical structures and category names visible in the images, or relevant clinical descriptions.
 $\bullet$ Learning ability.
 We begin by conducting an open-book test for each model to quantitatively measure how effectively each model has internalised the visual and linguistic patterns from the training phase.
-Specifically, we evaluate each model on the samples they have seen during training, i.e., 27,606 validation dialogues.
+Specifically, we evaluate each model on the samples they have seen during training, i.e., validation dialogues in ColonINST.
 The “seen” columns in Tab.
-VII show that we achieve the highest scores in the CLS ($\mathcal{A}\!=\!94.02\%$) and REC ($\text{IoU}\!=\!65.89\%$) tasks.
-This implies that ColonGPT has a better learning ability, which allows it to classify images and understand reference expressions related to specific visual regions.
-Furthermore, we achieve an accuracy of $99.02\%$, below $99.32\%$ achieved by LLaVA-v1.5 (#C2).
-This gap ($\Delta\!=\!0.3\%$) is due to a 7B-level LM used in LLaVA-v1.5, which has also been pretrained on additional data and thus gained more knowledge.
+VII show that we achieve the highest scores in the CLS ($\mathcal{A}\!=\!94.06\%$), REG ($\mathcal{A}\!=\!99.96\%$), and REC ($\text{IoU}\!=\!85.74\%$) tasks.
+This suggests that our ColonGPT has better learning ability, which allows it to correctly classify images and understand reference expressions related to specific visual regions.
 $\bullet$ Generalisation ability.
-We further conduct a closed-book test to examine each model’s ability to generalise knowledge to unseen conditions, i.e., 165,830 testing samples of ColonINST.
+We further conduct a closed-book test to examine each model’s ability to generalise knowledge to unseen conditions, i.e., testing samples of ColonINST.
 The “unseen” columns in Tab.
-VII consistently reveal our superior performance in unseen samples in all three tasks.
-Recall that our model performs worse than LLaVA-v1.5 in the REG task for seen scenarios.
-However, this gap is overtaken when exposed to unseen scenarios, where we achieve an accuracy of $83.42\%$, even exceeding the 7B-level LLaVA-v1.5 (#C2) by $10.54\%$.
-Compared to the runner-up model, LLaVA-Med-v1.5 (#H1), our model surpasses it by $1.2\%$ in the seen scenarios for the REC task, and this gap is further widened by $3.8\%$ in the unseen scenarios.
-These results show the potential of ColonGPT as a conversational assistant for multimodal colonoscopy tasks, especially in generalising to unseen data.
+VII consistently reveal our superior performance on unseen samples across all three tasks.
+Recall that our model slightly outperforms LLaVA-v1.5 in the REG task for the seen scenarios, i.e., $99.96\%$ (Ours) vs.
+$99.32\%$ (#C2).
+However, this gap is further widened when exposed to unseen scenarios, where we achieve an accuracy of $80.18\%$, even exceeding the 7B-level LLaVA-v1.5 (#C2) by $7.3\%$.
+Moreover, compared to medically tailored models, ColonGPT shows great potential as a colonoscopy-specific conversational assistant for multimodal tasks, especially in generalising to unseen data.
+For example, we exceed #G1 and #H1 models by $5.2\%$ and $4.0\%$, respectively in the unseen scenarios for the CLS task.
+Figure 7: Illustration of ColonGPT’s multimodal capabilities.
+Our model can execute various multimodal colonoscopy tasks through conversational interactions, including comprehension (CLS, REG), localisation (REC), and captioning (CAP) based.
 TABLE VIII: Diagnostic studies of three core components in ColonGPT.*”: interpolate the position embeddings for higher resolution, specifically from 224px to 384px.
@@
 Visual encoder input/URL CLS REG REC
-ConvNeXtV2-L 384px/link 82.95% 78.63% 33.74%
-ViT-L 384px/link 82.16% 77.04% 40.78%
-MAE-L* 384px/link 80.85% 75.87% 38.53%
-MAE-L 224px/link 81.95% 77.62% 43.25%
-DINOv2-L* 384px/link 35.03% 22.91% 6.79%
-DINOv2-L 224px/link 21.22% 7.96% 2.69%
-CLIP-L 336px/link 83.99% 78.67% 41.54%
-SigLIP-SO 384px/link 85.81% 83.42% 45.77%
+ConvNeXtV2-L 384px/link 81.51% 76.41% 53.09%
+ViT-L 384px/link 83.99% 78.79% 56.64%
+MAE-L* 384px/link 83.30% 79.51% 53.59%
+MAE-L 224px/link 82.85% 79.19% 57.57%
+DINOv2-L 224px/link 22.48% 10.09% 5.65%
+CLIP-L 336px/link 77.05% 71.92% 51.68%
+SigLIP-SO 384px/link 83.24% 80.18% 56.24%
 (b) Multigranuarity multimodal adapter
 token (ratio) CLS REG REC
-MLP baseline 729 (100.00%) 83.53% 81.80% 43.70%
-$\{16,8,1\}$ 321 (44.03%) 84.39% 80.90% 46.37%
-$\{14,7,1\}$ 246 (33.74%) 85.81% 83.42% 45.77%
-$\{14,7\}$ 245 (33.61%) 85.01% 82.49% 43.62%
-$\{12,6,1\}$ 181 (24.83%) 83.74% 81.60% 45.94%
-$\{10,5,1\}$ 126 (17.28%) 84.28% 82.01% 46.46%
-$\{8,4,1\}$ 81 (11.11%) 84.70% 81.36% 45.30%
-w/o Pos. Enc. 246 (33.74%) 84.50% 82.91% 40.09%
+MLP baseline 729 (100.00%) 78.10% 75.27% 56.29%
+$\{16,8,1\}$ 321 (44.03%) 79.84% 76.13% 53.15%
+$\{14,7,1\}$ 246 (33.74%) 83.24% 80.18% 56.24%
+$\{14,7\}$ 245 (33.61%) 80.67% 78.86% 58.05%
+$\{12,6,1\}$ 181 (24.83%) 79.46% 75.25% 52.22%
+$\{10,5,1\}$ 126 (17.28%) 81.42% 76.73% 52.04%
+$\{8,4,1\}$ 81 (11.11%) 81.55% 77.60% 53.39%
+w/o Pos. Enc. 246 (33.74%) 80.56% 76.35% 55.77%
 (c) Fine-tuning strategy
 Strategy $r$ $\alpha$ CLS REG REC
-full-tuning - - 78.06% 73.79% 50.20%
-LoRA 4 8 85.43% 82.75% 45.02%
-LoRA 8 16 84.45% 80.78% 44.98%
-LoRA 16 32 84.39% 80.81% 45.90%
-LoRA 32 64 84.91% 82.73% 45.56%
-LoRA 64 128 83.84% 81.19% 43.57%
-LoRA 128 256 85.81% 83.42% 45.77%
-LoRA 256 512 82.93% 79.96% 48.27%
+full-tuning - - 70.03% 63.07% 13.17%
+LoRA 8 16 80.51% 77.61% 51.61%
+LoRA 16 32 81.95% 77.87% 52.41%
+LoRA 32 64 82.63% 78.76% 52.11%
+LoRA 64 128 82.27% 79.72% 52.20%
+LoRA 128 256 83.24% 80.18% 56.24%
+LoRA 256 512 81.75% 77.89% 54.03%
+LoRA 512 1024 79.39% 74.54% 55.52%
 $\bullet$ Qualitative analysis.
@@
 Our model provides descriptions of a pedunculated polyp, detailing its morphology, contextual characteristics, and potential clinical relevance.
-Additionally, ColonGPT can describe the treatment procedure when an instrument is present, e.g., “The polyp is being examined using a snare.”
-5.3.2 Diagnostic studies
+Additionally, ColonGPT can describe the treatment procedure when an instrument is present, e.g., “The presence of a snare, a tool used to remove polyp”
+V-D Diagnostic experiments
 $\bullet$ Visual encoder.
@@
 As shown in Tab.
-(8-a), all encoders use pre-trained weights from Huggingface.
+(VIII-a), all encoders use pre-trained weights from Huggingface.
 To ensure consistency, we manually interpolate the smaller position embedding for MAE and DINOv2 from 224px to 384px (marked with *), while leaving the default input of the remaining models unchanged.
@@
 Regarding the other unimodal encoders, both supervised learning methods (ConvNeXtV2, ViT) and the reconstructive approach (MAE) give satisfactory feedback.
-However, the vision-only contrastive learning model (DINOv2) struggles, especially in the REC and REG tasks.
-Multiple efforts to optimise DINOv2 with various hyperparameters, including input size, model size, and learning rate, did not produce acceptable results for multimodal colonoscopy tasks.
+However, the vision-only contrastive learning model (DINOv2) struggles [^1] Notably, multiple efforts to optimise DINOv2 with various hyperparameters, including input size, model size, and learning rate, did not produce acceptable results for multimodal colonoscopy tasks., suggesting that its visual representations may be difficult to align with the language space.
 $\bullet$ Multigranularity multimodal adapter.
@@
 As detailed in Tab.
-(8-b), we analyse its effectiveness from three perspectives.
+(VIII-b), we analyse its effectiveness from three perspectives.
 (a) How to configure the pooling kernels?
@@
 It decreases the visual tokens from 100% (729 tokens) to 33.74% (246 tokens) while maintaining impressive results across three conversational tasks.
-To illustrate, we observe performance gains of 2.28%, 1.62%, and 2.07% in the CLS, REG, and REC tasks, respectively.
+To illustrate, we observe performance gains of 5.14% and 4.91% in the CLS and REG tasks, respectively.
 (b) Is global context necessary?
@@
 As shown in the last row of Tab.
-(8-b), our setup without positional encoding shows a significant performance drop in REC task, from 45.77% to 40.09%.
+(VIII-b), our model without positional encoding shows a significant performance drop in the REG task, from 80.18% to 76.35%.
 This suggests that the relative position information for the visual sequence is crucial for the localisation task.
@@
 As shown in Tab.
-(8-c), we initiate a set of variants to tune the language model, Phi1.5.
-It includes seven variants with different setups for LoRA and a full-tuning variant as a reference point.
-The best performance was observed in the LoRA variant with configuration $r/\alpha=128/256$, achieving 85.81% and 83.42% in the CLS and REG tasks, respectively.
-In addition, the other two variants achieve outstanding scores in the REC task, with the full-tuning variant achieving 50.20% and the LoRA with configuration $r/\alpha\!=\!256/512$ reaching 48.27%.
-This implies that more tunable parameters for the language model could allow ColonGPT to capture more complex patterns in the spatial planes.
-Therefore, we believe that there remains room to improve the localisation ability of ColonGPT in the REC task, such as increasing the rank factor $r$ and finding a matched scaling factor $\alpha$.
-For training efficiency, we finally choose $r/\alpha\!=\!128/256$ as our default configuration, as it allows us to reduce the training time from eight to roughly five hours.
-5.3.3 Empirical takeaways
+(VIII-c), we initiate a set of variants to tune the language model, Phi1.5.
+It includes seven variants with different LoRA setups and a full-tuning variant as a reference point.
+The best performance was observed in the LoRA variant with configuration $r/\alpha=128/256$.
+Moreover, the full-tuning variant yields unsatisfactory performance on three tasks.
+This suggests that an increase in the tunable parameters does not produce gains under the conditions of limited-scale data.
+V-E Empirical takeaways
 This study represents a preliminary exploration of multimodal instruction tuning techniques in colonoscopy.
@@
 This success benefits from the way we build ColonGPT.
-(a) Colonoscopy data inherently contains redundant information, such as the fact that most mucosal surfaces are similar, as well as camouflaged patterns between benign lesions and their surroundings, as discussed in §2.2.
+(a) Colonoscopy data inherently contains redundant information, such as the fact that most mucosal surfaces are similar, as well as camouflaged patterns between benign lesions and their surroundings, as discussed in §II-B.
 To reduce redundancy, we propose a multigranularity multimodal adapter that selectively samples tokens without compromising performance.
@@
 This indicates that larger models appear to require more colonoscopy data.
-Thus, future efforts should prioritise enhancing training and inference efficiency, especially for the medical field, rather than racing with massive computational resources.
+Thus, future efforts should prioritise enhancing parameter efficiency, especially for the medical field, rather than racing with massive computational resources.
 A promising idea to streamline the MLM framework using an encoder-free solution [293] to interpret visual pixels.
@@
 We encourage that the vision and language parts of the next-generation framework can handle their respective roles, such as a parallel framework [296] that predicts segmentation masks and generates language captions, simultaneously.
-6 Conclusion
+VI Conclusion
 We investigate the frontiers in intelligent colonoscopy techniques and their broader implications in the multimodal field.
@@
 Second, our survey reveals that multimodal research in colonoscopy is underexplored.
-To embrace this, we contribute three initiatives to the community: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-specific multimodal language model ColonGPT, and a multimodal benchmark.
-Importantly, multimodal colonoscopy research is rapidly advancing.
-Future developments can build on recent breakthroughs, such as executing role-playing tasks with agents [297].
+To this end, we contribute three initiatives to the community: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-specific multimodal language model ColonGPT, and a multimodal benchmark.
 Acknowledgments
-This research was supported by NSFC (NO.62476143). We express our sincere gratitude to Yu-Cheng Chou (JHU) and Stephen Gould (ANU) for interesting discussions.
+This work was supported by the ANU-Optus BRCoE (scholarship awarded to Ge-Peng Ji) and the NSFC (No.
+62476143).
+We express our sincere gratitude to Yu-Cheng Chou (JHU) and Stephen Gould (ANU) for their interesting discussions.